### v3

[…]
Hagn et al.
[…]
Stochastic Team Formation and Routing
[…]
A Branch-Price-Cut-And-Switch Approach for Optimizing Team Formation and Routing for Airport Baggage Handling Tasks with Stochastic Travel Times
[…]
AdONE GRK 2201
[…]
AdONE GRK 2201
[…]
Department of Discrete Mathematics
[…]
In airport operations, optimally using dedicated personnel for baggage handling tasks plays a crucial role in the design of resource-efficient processes. Teams of workers with different qualifications must be formed, and loading or unloading tasks must be assigned to them. Each task has a time window within which it can be started and should be finished. Violating these temporal restrictions incurs severe financial penalties for the operator. In practice, various components of this process are subject to uncertainties. We consider the aforementioned problem under the assumption of stochastic travel times across the apron. We present two binary program formulations to model the problem at hand and propose a novel solution approach that we call Branch-Price-Cut-and-Switch, in which we dynamically switch between two master problem formulations. Furthermore, we use an exact separation method to identify violated rank-1 Chvátal-Gomory cuts and utilize an efficient branching rule relying on task finish times. We test the algorithm on instances generated based on real-world data from a major European hub airport with a planning horizon of up to two hours, 30 flights per hour, and three available task execution modes to choose from. Our results indicate that our algorithm is able to significantly outperform existing solution approaches. Moreover, an explicit consideration of stochastic travel times allows for solutions that utilize the available workforce more efficiently, while simultaneously guaranteeing a stable service level for the baggage handling operator.
[…]
airport operations, team formation, routing, hierarchical skills, uncertainty, branch-price-and-cut
1 Introduction
Baggage handling tasks, namely loading and unloading containers and bulk luggage, play an important role in aviation and airport operations. They are one of several ground handling tasks (Evler et al. 2021). Whenever a plane is arriving or scheduled for departure, a team of workers that operates several types of equipment needs to be formed and assigned to load or unload the aircraft. Said workers have different skill levels, i.e., qualifications for the types of equipment, such as tractors or high cargo loaders, they are allowed to operate. Depending on the airplane model and type, various equipment compositions can be used to execute the (un-)loading task, leading to different workforce requirements. For instance, large-sized airplanes typically have at least two cargo holes. These can be unloaded either sequentially or in parallel. For a more detailed description of the processes, see Dall’Olio and Kolisch (2023). In the following, whenever loading tasks are discussed, both loading and unloading processes are encompassed.
Each type of equipment requires a certain skill level, where skill levels are ordered hierarchically. That is, a worker with skill level $k$ can execute any job requiring a skill level of $k$ or lower. Frequent delays in baggage claim processes and plane departures can result from improper team formation and task assignment decisions. This noticeably decreases passenger satisfaction and induces significant financial penalties for the baggage handling operator. Eurocontrol (2023) state that around 30% of all aircraft delays are caused by airline-related issues, to which baggage handling processes are attributed. Furthermore, depending on the airport and the considered planning horizon, baggage handling is subject to significant uncertainties regarding travel times of ground vehicles. Delays while traveling from one parking position to another, e.g., because of planes crossing the apron, are quite frequent and typically accumulate to a significant scope. In the following, we consider loading times to be deterministic, while travel times are assumed to be stochastic with known probability distributions. Moreover, to limit the potential financial penalties for the baggage handling operator, it is reasonable to demand that each task’s time window is satisfied with at least a predefined probability. Our work builds upon Dall’Olio and Kolisch (2023), which addresses the deterministic team formation and routing problem. The main contributions of this paper are:
  • i.
    We extend previous works on baggage handling optimization by including stochastic travel times.
  • ii.
    We propose a novel Branch-Price-Cut-and-Switch solution approach that dynamically switches between two master problem formulations, depending on the solution’s characteristics.
  • iii.
    We conduct extensive experimental studies to analyze the impact of stochastic information on optimal solutions and assess our solution method’s efficiency.
The remainder of this paper is structured as follows. Section 2 reviews the most relevant literature for our study. Section 3 provides a detailed problem description with the help of mathematical notation. In Section 4, we propose two binary programs that are used to model the problem at hand. In Section 5, we develop a Branch-Price-Cut-and-Switch solution approach and elaborate on its core components. Section 6 presents computational experiments aiming at assessing the proposed algorithm’s performance and the impact of stochasticity on optimal solutions. We summarize our findings in Section 7 and present several areas of future research.
2 Literature Review
The problem considered in this paper can be seen as a variant of the technician scheduling and routing problem (TSRP). TSRPs typically consist of a routing part that can be seen as a vehicle routing problem (VRP) and a scheduling or team formation part. Section 2.1 focuses on literature dealing with stochastic VRPs in a general context. Section 2.2 considers publications dealing with stochastic TSRP variants.
2.1 Literature on Stochastic Vehicle Routing
Recently, stochastic formulations of the vehicle routing problem with time windows (VRPTW) have seen a noticeable interest increase. In the following, we provide an overview of literature dealing with stochastic variants of the VRPTW. Oyola, Arntzen, and Woodruff (2018) summarize the most relevant literature on common types of stochastic VRPs. A taxonomy and overview of the space of VRPs can be found in Eksioglu, Vural, and Reisman (2009).
Generally, random variables in the stochastic VRPTW are assumed to be independent as stochastic dependency drastically complicates the calculation of distributions and their moments. While some publications focus on the VRPTW with stochastic demands (see Lee, Lee, and Park (2012), Zhang, Lam, and Chen (2016)), stochastic travel and service times are far more common. Stochasticity can be addressed in different ways. A common approach is to include chance constraints, which limit the probability of violating time windows. While Errico et al. (2018) connect chance constraints to the simultaneous satisfaction of all time windows, Ehmke, Campbell, and Urban (2015) as well as Li, Tian, and Leung (2010) limit the probability of violating the time window of each customer individually.
In vehicle routing problems, the distributions of start times at a given customer depend on the sequence of customers previously visited by the same vehicle and the distributions of travel and service times along the route. Furthermore, the computational effort required to calculate said distributions heavily depends on the type of time windows. Taş et al. (2014) assume that time windows are soft and travel times are independent, gamma distributed random variables. The authors propose an exact calculation of arrival time distributions by convolving finish time and travel time distributions and solve the problem using a Branch-Price-and-Cut approach. Because the convolution of gamma distributed random variables is gamma distributed, finish time distributions are again gamma distributed with known shape and scale parameter, which significantly simplifies their calculation.
If time windows are assumed to be hard, arrival time distributions need to be truncated at the start of each time window. Hence, distributional properties, such as moments or the existence of a closed-form expression, are not propagated along vehicle routes. Therefore, distributions can only be obtained by explicitly calculating the convolution of two distributions of arbitrary shapes, which is often not computationally tractable in reasonable time. Errico et al. (2018) assume service times to be stochastic, discretely distributed random variables and calculate start time distributions exactly. They use a Branch-Price-and-Cut algorithm to solve the problem and report computational results for four different variants of triangular distributions. Miranda and Conceição (2016) assume normally distributed service and travel times. They approximate service and travel time distributions by discretizing them on a finite set of values, whose count and values depend on the underlying distribution’s moments. This allows for an approximation of the finish time distribution of a given task by convolving the finish time distribution of the previous task with the aforementioned approximations of travel and service time distributions. They further improve their approach in Miranda, Branke, and Conceição (2018) using improved lower bounds for the approximation intervals of service and travel times. Ehmke, Campbell, and Urban (2015) show that the first and second moment of start time distributions can be calculated with little effort if both travel times and start times at the previous customer are normally distributed. While the former is usually not the case, the authors present computational results that indicate that their approach works well for normal, shifted gamma, and shifted exponential travel time distributions. Li, Tian, and Leung (2010) use stochastic simulation to derive estimates for start time distributions and solve the problem using tabu search.
Our considerations can be seen as a combination and extension of Errico et al. (2018) and Li, Tian, and Leung (2010), as we calculate start time distributions exactly but interpret chance constraints as a non-route-interdependent property and consider a stochastic objective function.
2.2 Literature on the stochastic TSRP
Unlike for the vehicle routing problem, there is no standard definition for the TSRP. In general, the TSRP consists of scheduling workers or assembling teams of workers and routing them across available tasks such that each task is executed by exactly one team (or worker). Depending on the context, additional requirements must be considered.
We first focus on literature on deterministic variants of the TSRP that share several key properties with the problem at hand. Pereira, Alves, and de Oliveira Moreira (2020) consider a multiperiod workforce scheduling and routing problem, where tasks do not have a fixed time window. However, precedence relationships need to be satisfied. The set of available teams and their characteristics, such as skill levels, are fixed and thus are not part of the decision space. The authors propose a mixed-integer model, which is shown to be computationally tractable only for small instances, and an ant-colony optimization heuristic. Çakırgil, Yücel, and Kuyzu (2020) examine a multi-objective workforce scheduling and routing problem, where teams consisting of workers with different skills need to be formed and task sequences need to be assigned to each team. Tasks can be started at any time, but need to be finished before their respective deadline. Workers can have different qualifications, but it is not possible for them to execute tasks that require a different qualification, i.e., downgrading is not possible. The authors developed a mixed-integer program and a 2-stage matheuristic to solve the problem. While the former proves to be inefficient for large instances, the latter is able to scale well with instance size. Li, Lim, and Rodrigues (2005) consider a manpower allocation problem with hard time windows. Similar to Çakırgil, Yücel, and Kuyzu (2020), workers with different qualifications need to be grouped into teams, which then execute a to-be-optimized sequence of tasks. Additionally, downgrading workers is not possible. The objective is to minimize a weighted sum of the total number of required workers and the total travel time. The authors develop a construction heuristic and a simulated annealing approach to solve the problem.
All of the publications previously discussed differ from our considerations in two aspects. First, they fix teams beforehand or make restrictive assumptions regarding team formation possibilities, such as the impossibility of downgrading or the existence of a single mode. Second, they focus on heuristic solution methods to solve medium- and large-sized instances. For an overview of further literature dealing with deterministic variants of the TSRP relevant to this study, we refer to Dall’Olio and Kolisch (2023).
While there is plenty of literature regarding the deterministic TSRP, very little research has been done on stochastic problem variants. In the following, we turn our focus on literature dealing with stochastic variants of the TSRP that are similar to ours.
Souyris et al. (2013) propose a robust formulation for the TSRP. Workers are assumed to be homogeneous and service times are stochastic. Each task must be started before a fixed deadline with a given probability. Moreover, each task must be executed by a single technician, thus team formation is not part of the model. The authors propose three different bounded uncertainty sets that limit the total service time delays per technician or client, respectively. The objective is to minimize the worst-case total delay and travel time. A branch-and-price approach is proposed and applied to real-world instances with 41 customers and 15 technicians.
Yuan, Liu, and Jiang (2015) address the TSRP in the context of health care workers and home health care services. Each task must be executed by a single worker. Furthermore, it is possible to downgrade workers to lower skill levels. Service times are assumed to be stochastic with a known probability distribution. The goal is to minimize the expected total travel costs, fixed costs of caregivers, service costs, and penalties for late arrival at customers. The authors use a branch-and-price approach to solve the problem exactly. Computational experiments assuming uniformly distributed service times and 25 to 50 customers indicate that the approach is able to provide very good results for small-sized instances. While the objective function and interpretation of stochasticity considered by Yuan, Liu, and Jiang (2015) is different from ours and several properties, such as the possibility of executing tasks using different modes and skill compositions, are missing, their solution approach exhibits similarities with the one proposed in the following and can be applied to the problem at hand. In Section 6.2, we compare both methodologies on a set of test instances.
Binart et al. (2016)
consider a TSRP with mandatory and optional tasks, where time windows are hard and fixed, and predefined teams must be used to execute tasks. Although not explicitly done, their modeling framework would allow considering multiple skill levels, downgrading, and multiple modes for single tasks. They assume that travel and service times are discretely distributed according to triangular distributions. Furthermore, the time windows of each task must be satisfied with a given probability. A 2-stage approach is proposed to heuristically solve the problem. First, a feasible skeleton solution is obtained using a generic MIP solver by optimally covering all mandatory tasks. Second, the first-stage solution is refined by inserting optional customers into the existing routes such that time window restrictions are not violated. Computational results on instances with 5 to 9 mandatory and 30 to 50 optional tasks indicate that the approach can yield good results in most cases. The problem considered by Binart et al. (2016) can be seen as a special case of the problem at hand where all technicians have the same skill level, each team formation represents one technician, and each task can be executed by each team formation. While the objective function and interpretation of stochasticity differ from ours, the models and solution approach proposed in the following can be applied to their problem.
Our work shows several similarities with Binart et al. (2016) and Yuan, Liu, and Jiang (2015), such as the stochasticity of travel times, the usage of exact solution methods, and the incorporation of several team properties such as multiple skill levels and modes. At the same time, our interpretation of service levels at individual tasks rather than entire solutions, the lack of assumptions on distributional properties, and the exact calculation of arrival times separate our considerations from previous approaches.
To conclude this section, it is worth noting that existing research primarily deals with heuristic solution procedures. Exact approaches are typically only efficient for small-sized instances, especially when stochasticity is considered.
The following considerations are widely based on Dall’Olio and Kolisch (2023), which consider the problem of forming and routing worker teams for baggage handling tasks in a deterministic setting. We extend their deterministic problem description by assuming stochastic travel times. Furthermore, we advance their proposed solution algorithm by introducing cutting planes and additional branching strategies, improving the runtime efficiency of the pricing step, and describing an efficient way to handle infeasible solutions that is based on dynamically switching between two master problem formulations.
3 Problem Description
In the following section, we provide a detailed description of the problem considered in this publication with the help of mathematical notation, which will be used throughout the following sections. Table 1 provides an overview of general notation introduced in this section and used throughout the following sections.
Table 1: General notation
Sets and variables
$\mathcal{Q}$ Set of profiles
$\mathcal{Q}_{i}$ Set of profiles compatible with task $i$
$\mathrm{Q}^{\mathrm{D}}$ Set of disaggregated profiles $(q,s)$
$\mathcal{S}$ Set of skill compositions
$\mathcal{S}_{q}$ Set of skill compositions compatible with profile $q$
$\mathcal{I}$ Set of tasks
$\mathcal{I}^{r}$ Set of tasks executed by route $r$
$\mathcal{K}$ Set of worker skill levels
$\mathcal{T}$ Set of discrete time points
$B_{i,j}$ (finite) support of travel time delays for tasks $i$ and $j$ (or tasks and the depot)
$\Omega$ Set of all possible travel time realizations
$t$ random variable of stochastic travel times
$t_{i,j}$ random variable of stochastic travel times between nodes $i$ and $j$
Input Parameters
$N_{k}^{\mathrm{D}}$ Number of available workers with skill level $k$
$N_{k}$ Number of available workers with at least level $k$
$\xi_{q,k}$ Number of workers with at least level $k$ required by working profile $q$
$d$ Central worker depot
$\mathrm{ES}_{i}$ Earliest start time of task $i$
$\mathrm{LF}_{i}$ Latest finish time of task $i$
$\mathrm{LF}_{i}^{\mathrm{e}}$ Extended latest finish time of task $i$
$w_{i}$ Weight of task $i$
$p_{i,q}$ Processing time of task $i$ when undertaken with working profile $q$
$\alpha$ Chance constraint probability (minimum service level requirement)
$t^{\mathrm{det}}$ Vector of deterministic, best-case travel times
$\omega$ (single) realization of travel times
$\omega_{\mathrm{max}}$ worst-case realization of travel times
Model Parameters
$P_{i}(\cdot)$ Penalty function for time window violations at task $i$
$\mathbb{E}(c^{r})$ Expected cost of route $r$
$\mathrm{tl}^{r,q}$ Leave time from the depot for the team with profile $q$ executing route $r$
$\mathrm{tr}^{r,q}$ Return time to the depot for the team with profile $q$ executing route $r$
$S^{r,q}_{i}$ (Stochastic) start time of task $i$ in tour $r$ with profile $q$
$F^{r,q}_{i}$ (Stochastic) finish time of task $i$ in tour $r$ with profile $q$
$F^{r,q}_{i}(\omega_{\mathrm{max}})$ Worst-case finish time of task $i$ in tour $r$ with profile $q$
General Notation
We consider a set $\mathcal{I}=\{1,\dots,|\mathcal{I}|\}$ of tasks that have to be performed within a specified planning horizon. Each task corresponds to either a loading or unloading task of an incoming (or outgoing) flight. We note that, in the context of VRPs, tasks are typically referred to as customers or customer visits. A time window $\mathrm{ES}_{i},\mathrm{LF}_{i}]$ is associated with each task $i\in\mathcal{I}$, describing the earliest and latest point at which service on an airplane can be started and should be finished, respectively.
Initially, a workforce of prespecified size is located at a central depot denoted by $d$. In general, workers can be subdivided into skill levels $\mathcal{K}\coloneqq\{1,\dots,K\}$. Depending on their skill level, workers are only allowed to operate certain equipment. Skill levels are ordered hierarchically, i.e., a worker with skill level $k$ is able to operate any equipment requiring level $l\leq k$. Moreover, the amount of workers with skill level $\geq k$ that are available at the worker depot is fixed and defined as $N_{k}\coloneqq\underset{l=k}{\overset{K}{\sum}}N_{l}^{\mathrm{D}}$, where $N_{k}^{\mathrm{D}}\in\mathbb{N}_{>0}$ is the number of available workers with skill level $k$. We note that in practical instances, higher-skilled workers tend to be rather scarce and hard to obtain as they require lengthy and expensive training. A route is a sequence $r\coloneqq\left(d,v_{1},\dots,v_{l},d\right)$ with $l\geq 1$ and $v_{i}\in\mathcal{I}$ for all $i=1,\dots,l$. Each team route $(r,q)$ has time instants $\mathrm{tl}^{r,q}$ and $\mathrm{tr}^{r,q}$ at which the team leaves the depot and returns to it, respectively. The set of tasks executed by route $r$ is denoted by $\mathcal{I}^{r}\coloneqq\{v_{1},\dots,v_{l}\}$. Furthermore, each route $r$ is assigned a profile $q=\left(\xi_{q,k}\right)_{k\in\mathcal{K}}\in\mathcal{Q}$ which is used to execute the route, where $\xi_{q,k}\in\mathbb{N}_{\geq 0}$ denotes the number of required workers of skill level greater than or equal to $k$ and $\mathcal{Q}$ is the set of available profiles. Because each type of aircraft requires different prerequisites and equipment for baggage handling tasks, only a subset $\mathcal{Q}_{i}\subseteq\mathcal{Q}$ of profiles can be used to execute a given task $i\in\mathcal{I}$. Depending on the profile $q\in\mathcal{Q}_{i}$, executing task $i$ takes $p_{i,q}\in\mathbb{N}_{>0}$ units of time.
Profiles and Skill Compositions
By construction, a profile $q\in\mathcal{Q}$ is well-defined by its aggregated workforce requirements, i.e., lower bounds for the skill levels of required workers, but does not contain any information about their actual skill levels. To include this information, we define a skill composition as a vector $s\in\mathbb{N}^{|\mathcal{K}|}$ and say that a skill composition $s=(s_{q,k})_{k\in\mathcal{K}}$ be can assigned to a profile $q\in\mathcal{Q}$ if
$\xi_{q,k^{*}}$ $=\underset{k\in\mathcal{K}}{\sum}s_{q,k}\quad\text{and}$ (1)
$\xi_{q,k^{\prime}}$ $\leq\underset{k\in\mathcal{K}:k\geq k^{\prime}}{\sum}s_{q,k}\quad\forall k^{\prime}\in\mathcal{K}$ (2)
hold, where $k^{*}=1$ is the skill level with the lowest qualification. Furthermore, the set
[…]
consists of all skill compositions that can be assigned to profile $q\in\mathcal{Q}$. Hence, we define the set
$\mathcal{Q}^{\mathrm{D}}\coloneqq\left\{(q,s):q\in\mathcal{Q},s\in\mathcal{S}_{q}\right\}.$
as the set of disaggregated profiles, i.e., profiles enhanced by information on the skill levels of workers used to assemble a team with profile $q$. Given a route $r$, we define the set of profiles that can be used to execute $r$ as $\mathcal{Q}^{r}\coloneqq\underset{i\in\mathcal{I}^{r}}{\bigcap}\mathcal{Q}_{i}$. An example illustrating the relationship between profiles and skill compositions is presented in Section 4.2.
Travel Times
After finishing a task $i$, a team with profile $q$ can either return to the depot $d$ and become available for regrouping or directly continue with another task $j$. Traveling between locations $i,j\in\mathcal{I}\cup\left\{d\right\}$, where $i$ and $j$ are either two tasks or one task and the depot, takes at least $t_{i,j}^{\mathrm{det}}\in\mathbb{N}_{\geq 0}$ units of time and does not depend on the profile. We assume travel times to be symmetric and time-independent. In practice, these travel times are subject to delays caused by unexpected, exogenous factors, such as aircrafts crossing the apron or local traffic congestion. In such situations, a common approach is to solve a multi-stage stochastic problem, where an initial solution is constructed and, once more information on stochastic parameters such as travel times becomes available over time, is readjusted if needed. However, as ground vehicle traffic is rather unpredictable and an aircraft’s movements are typically only determined moments before the plane starts moving, it is not possible to avoid the usage of congested parts of the apron by re-adjusting the previously derived solution, i.e., by changing the sequences of tasks that should be executed by the same team or assigning a new team to a given task. To formalize this, we define non-negative, finite supports for travel time delays for each pair $(i,j)$ of tasks (or tasks and the depot) by
$B_{i,j}\subset\mathbb{N}_{\geq 0},\ |B_{i,j}|<\infty\quad\forall(i,j)\in E$
where $E\coloneqq\left\{(i,j):i,j\in\mathcal{I}\cup\{d\},\ i\neq j\right\}$. We then define the set of possible travel time delays by
$\Omega\coloneqq\underset{(i,j)\in E}{\bigtimes}B_{i,j}.$
Then, the vector of stochastic travel times $t:\ \Omega\rightarrow\mathbb{N}_{>0}^{|E|}$ is given by
$t(\omega)$ $=t^{\mathrm{det}}+\omega\quad\forall\omega\in\Omega$
for all $(i,j)\in E$. Consistently, we can represent $t(\omega)$ as $t(\omega)=(t_{i,j}(\omega))_{(i,j)\in E}=(t_{i,j}^{\mathrm{det}}+\omega_{i,j})_{(i,j)\in E}$. We assume that events in $\Omega$ are independent, thus $(t_{i,j})_{(i,j)\in E}$ are independent, non-negative random variables with finite support and a known probability distribution $\Xi_{i,j}:\mathbb{N}_{\geq 0}\rightarrow[0,1]$ with $\Xi_{i,j}(t)\coloneqq\mathbb{P}(t^{det}_{i,j}+\omega=t)$ and $\underset{t\in\mathbb{N}_{\geq 0}}{\sum}\Xi_{i,j}(t)=1$. In the following, $t_{i,j}^{\mathrm{det}}$ is also called deterministic or best-case travel time. Additionally, we denote the worst-case scenario of travel times by
$\omega_{\mathrm{max}}=\left(\max\{\omega_{i,j}:\omega_{i,j}\in B_{i,j}\}\right)_{(i,j)\in E}$
Routes and Route Feasibility
If a worker team arrives at task $i$ before its time window opens, it has to wait until $\mathrm{ES}_{i}$ to start the service. Because travel times are stochastic, the start and finish times are random variables, which depend on the worker profile $q$ used to execute the route, the sequence $r$ of tasks, and a realization $\omega\in\Omega$ of travel times. Formally, we define the start and finish times $S_{i}^{r,q}:\Omega\rightarrow\mathbb{N}$ and $F_{i}^{r,q}:\Omega\rightarrow\mathbb{N}$ as
$S_{i}^{r,q}(\omega)$ $\coloneqq\max\{F_{i-1}^{r,q}(\omega)+t_{i-1,i}(\omega),\mathrm{ES}_{i}\},\quad\text{and}$
$F_{i}^{r,q}(\omega)$ $\coloneqq S_{i}^{r,q}(\omega)+p_{i,q},$
where $F_{d}^{r,q}(\omega)=\mathrm{tl}^{r,q}$ is equal to the depot leave time of team route $(r,q)$ and i1i-1 is the predecessor of task $i$ along $r$. Furthermore, we denote by $\mathrm{tr}^{r,q}$ the worst-case return time of team route $(r,q)$, which is realized when the worst-case scenario $\omega_{\mathrm{max}}$ occurs.
As the actual travel times are random and hardly predictable beforehand, finishing tasks after the time window closes, i.e., after $\mathrm{LF}_{i}$, is sometimes inevitable. Nevertheless, baggage handling operators often have to guarantee a minimum service level, i.e., the percentage of flights that are not delayed due to improper baggage handling. Moreover, in order to limit the probability of passengers missing their connecting flights, airlines often specify maximum acceptable delays in case of disruptions. Therefore, we allow tasks to be finished after their time window closes, while we limit said violation to at most a fixed amount of time, leading to a new extended latest finish time $\mathrm{LF}_{i}^{\mathrm{e}}\geq\mathrm{LF}_{i}$ for each task $i$. Furthermore, to limit potential financial damages and customer dissatisfaction, we introduce a minimum service level $\alpha\in[0,1]$ and demand that the finish time $F_{i}^{r,q}$ of each task $i$ on each route $r$ must satisfy
$\mathbb{P}(F_{i}^{r,q}\leq\mathrm{LF}_{i})\geq\alpha,$ (3)
$\mathbb{P}(F_{i}^{r,q}>\mathrm{LF}_{i}^{\mathrm{e}})=0.$ (4)
Constraint (3), also called chance constraint or service level constraint, limits the probability of delays caused by the baggage handling operator, while constraint (4) guarantees that potential delays do not exceed a prespecified limit. A route $r$ with profile $q$ is called feasible if (3)–(4) are satisfied for all $i\in\mathcal{I}^{r}$. We note that, for planning horizons of up to a few hours, the arrival times of aircraft are typically subject to very little uncertainty as there are no major causes for delay once a plane has departed from its origin airport. Moreover, delays of departing flights, i.e., loading tasks, are either known far in advance or realize shortly before the time window opens. In the former case, delays can be incorporated into the definition of time windows. In the latter case, one is typically unable to readjust the current plan as the loading has already begun and, if said delays were modeled as stochastic time windows, one might obtain overly robust solutions due to the low probability of such events happening. Therefore, the assumption of deterministic time windows is reasonable for short- to medium-term planning horizons.
Route Costs and Objective Function
Each task $i\in\mathcal{I}$ has an assigned weight $w_{i}\geq 0$ that indicates its importance in the flight schedule. This can be, for instance, the number or percentage of passengers that have to reach a connecting flight at the destination airport. The expected cost of a route $r$ with profile $q$ is then given by
$\mathbb{E}(c^{r})=\underset{i\in\mathcal{I}^{r}}{\sum}w_{i}\cdot\mathbb{E}\left(\left(F_{i}^{r,q}-\mathrm{EF}_{i}\right)+P_{i}(F_{i}^{r,q})\right)$
where $\mathrm{EF}_{i}\coloneqq\mathrm{ES}_{i}+\min\{p_{i,q}:\ q\in\mathcal{Q}_{i}\}$ is the earliest possible finish time of task $i$ and $P_{i}:(\mathrm{LF}_{i},\mathrm{LF}_{i}^{\mathrm{e}}]\rightarrow\mathbb{R}_{>0}$ is a penalty function. The first part of the objective functions aims to minimize weighted expected finish times in order to have as large safety time buffers as possible to absorb delays in other parts of the baggage handling process. The second part consists of a penalty for delaying flights. Because the probability of passengers missing connecting flights increases superlinearly in the delay of a given baggage handling task, the baggage handling operator aims to keep the variance of delays as low as possible. As the variance can be seen as a quadratic function of the delay, we use a quadratic penalty function
$P_{i}(F_{i}^{r,q}(\omega))\coloneqq\begin{cases}\begin{aligned} &(F_{i}^{r,q}(\omega)-\mathrm{LF}_{i})^{2}&&\text{if}\ \ F_{i}^{r,q}(\omega)>\mathrm{LF}_{i},\\ &0&&\text{otherwise}\end{aligned}\end{cases}$
for all $\omega\in\Omega$. In order to calculate the expected cost of a route $r=(d,v_{1},\dots,v_{l},d)$ with $l\geq 1$, it is necessary to have full information about the finish time distribution of each task assigned to $r$. Let $i,j\in\mathcal{I}$ be two tasks that are executed consecutively in route $r$, i.e., there exists an $h\in\{1,\dots,l-1\}$ with $v_{h}=i$ and $v_{h+1}=j$. If the distribution of $F_{i}^{r,q}$ is known, the start time distribution of the subsequent task $S_{j}^{r,q}$ can be calculated as
$\mathbb{P}(S^{r,q}_{j}=\tau)=\begin{cases}\begin{aligned} &\underset{z=0}{\overset{\mathrm{ES}_{j}}{\sum}}\mathbb{P}(F^{r,q}_{i}+t_{i,j}=z)&&\text{if}\ \ \tau=\mathrm{ES}_{j}\\ &\mathbb{P}(F^{r,q}_{i}+t_{i,j}=\tau)&&\text{if}\ \ \tau>\mathrm{ES}_{j}\\ &0&&\text{otherwise}\end{aligned}\end{cases}$ (5)
The finish time distribution $F_{j}^{r,q}$ can then be obtained because $F_{j}^{r,q}(\omega)=S_{j}^{r,q}(\omega)+p_{j,q}$ holds by definition for all $\omega\in\Omega$.
4 Two Binary Program Formulations
In the following sections, we present two set-covering formulations for the problem. In Section 4.1, we extend the model proposed in Dall’Olio and Kolisch (2023) to incorporate stochastic travel times and a stochastic objective function. Besides the construction of routes, said model only decides on the profiles of teams, i.e., on lower bounds for the actual qualifications of required workers. Therefore, skill levels are implicitly aggregated, i.e., they are only considered on an aggregate level. Hence, one might obtain an integer solution that is operationally infeasible, i.e., it can not be implemented in practice. Whenever such a solution is found, we switch to an alternative formulation that considers workers based on individual skill levels. This alternative formulation, which always returns operationally feasible solutions but is considerably harder to solve, is presented in Section 4.2, alongside additional insights into operational feasibility.
4.1 Master Problem with Aggregated Workforces
The following considerations assume an underlying finite time grid given by discrete time points $\tau\in\mathcal{T}$. We first introduce general notation and several auxiliary parameters. We define a team route as a tuple $(r,q)$ consisting of a feasible route $r$ and an associated worker profile $q\in\mathcal{Q}^{r}$, which is used to execute the route. The set of team routes is denoted by $\mathcal{R}$.
Let $(r,q)\in\mathcal{R}$ be a team route. We define by $b_{k,\tau}^{r,q}(\omega_{\mathrm{max}})\in\mathbb{N}_{\geq 0}$ the number of workers of skill level $\leq k$ that are conducting route $r$ and are occupied at time $\tau$, given worst-case travel times $t(\omega_{\mathrm{max}})$. This value is equal to $\xi_{q,k}$ for all $\tau\in[\mathrm{tl}^{r,q},\mathrm{tr}^{r,q}]$ and 0 else. Binary variables $\lambda^{r}_{q}$ indicate if team route $(r,q)$ is part of the solution. The problem can then be described using the following aggregated master problem, short AMP:
$\min$ $\sum_{(r,q)\in\mathcal{R}}\mathbb{E}(c^{r})\lambda^{r}_{q}$ (6)
s.t. $\underset{(r,q)\in\mathcal{R}:\ i\in\mathcal{I}^{r}}{\sum}\lambda^{r}_{q}\geq 1\quad\forall i\in\mathcal{I}$ (7)
$\underset{(r,q)\in\mathcal{R}}{\sum}b_{k,\tau}^{r,q}(\omega_{\mathrm{max}})\lambda^{r}_{q}\leq N_{k}\quad\forall k\in\mathcal{K},\ \forall\tau\in\mathcal{T}$ (8)
$\lambda^{r}_{q}\in\mathbb{N}\quad\forall(r,q)\in\mathcal{R}$ (9)
Constraints (7) ensure that each task is part of at least one route. It is easy to see that covering a task more than once can not be part of an optimal solution. Inequalities (8) guarantee that the workforce required at any given time $\tau$ does not exceed the total available workforce on an aggregate level. The objective function (6) aims to minimize the expected finish times of tasks and incurred penalties in order to maximize buffer times. Furthermore, the AMP with only a subset of team routes $\bar{\mathcal{R}}\subset\mathcal{R}$ considered and integrality conditions (9) relaxed to $\lambda^{r}_{q}\geq 0$ is called aggregated reduced master problem (ARMP). However, for the sake of simplicity, we will still be referring to the set of columns of the ARMP as $\mathcal{R}$. We note that, unlike in standard set covering formulations, we only require the variables to be integral rather than binary. However, for any feasible instance of the AMP, there exists an optimal solution which is binary. A proof for this can be found in Appendix 9.
We note that chance constraints (3) are not route-interdependent, thus they can be fully embedded in the pricing problem.
4.2 Master Problem with Skill-Level Specific Workforces
In the following, we provide a more detailed master problem that considers the workforce on an individual skill level basis. Furthermore, we provide some insights into the dominance relation between the two proposed master problem formulations.
Let $(r,q)\in\mathcal{R}$ be a team route. As described in Section 3, $q$ is well-defined by the amount $\xi_{q,k}$ of required workers of skill level greater or equal than $k$ for all $k\in\mathcal{K}$. This information suffices to ensure that the available workforce is never exceeded on an aggregate level, but does not guarantee feasibility in terms of allocation of workers to individual tasks.
Example 4.1
Consider the following example by Dall’Olio and Kolisch (2023). There are two workers $w_{1}$ and $w_{2}$ with a skill level of $1$ and $2$, respectively, available. Assume the optimal solution consists of four tours $T_{1},T_{2},T_{3}$, and $T_{4}$ depicted in Figure 1. Tours $T_{1}$ and $T_{4}$ each require one worker with skill level $2$ while tours $T_{2}$ and $T_{3}$ each require one worker with skill level $1$. It is clear to see that the solution is feasible for the AMP. Furthermore, worker $w_{2}$ must be assigned to tours $T_{1}$ and $T_{4}$, as these require a worker with skill level $2$. Worker $w_{1}$, however, cannot be assigned to both tours $T_{2}$ and $T_{3}$ as these overlap. Therefore, while the solution selects feasible team formations for all required teams, i.e., workforce constraints are satisfied on an aggregated level, there is no feasible selection of skill compositions that does not violate workforce restrictions. Hence, the team formation, as implied by the given solution, is not implementable in practice.
Figure 1: Tour Schedule satisfying (7)–(8) (from Dall’Olio and Kolisch (2023))
To deal with such undesirable solutions, the aforementioned authors propose an integer problem called feasibility check, which tries to find a feasible allocation of available workers to teams and regrouping strategies by solving a network flow problem on an appropriate graph. If said problem is infeasible, a cut is added to the master problem and the ARMP is re-solved. In preliminary studies, we observed that there are several instances in which a large number of cuts have to be added, degrading the algorithm’s performance by a large margin. In such instances, there is a large cardinality of binary solutions to the ARMP that all share the property of being operationally infeasible in the previously described sense and have the same (or almost the same) objective function value. This makes it necessary to consecutively forbid these solutions one by one. Furthermore, each additional cut introduces a dual value into the pricing problem, whose impact on the reduced cost of a column is only observable after a column has been generated. Therefore, it is possible that during the pricing procedure, columns that are part of a forbidden solution are re-generated. Thus, at the end of each pricing step, such columns need to be identified and discarded, increasing the computational effort required to solve the pricing problem. Additionally, the ARMP has to be re-solved every time without improving the solution quality.
We mitigate this issue by utilizing the concept of skill compositions to develop an alternative formulation to the AMP. For this, we define a disaggregated team route $(r,q,s)$ as a tuple consisting of a team route $(r,q)\in\mathcal{R}$ and a disaggregated profile $(q,s)\in\mathcal{Q}^{\mathrm{D}}$ and denote the set of disaggregated team routes by $\mathcal{R}^{\mathrm{D}}$.
Example 4.2
Consider the following example of a team route $(r,q)\in\mathcal{R}$ with $r\in\mathcal{R}$, $q\in\mathcal{Q}^{r}$ . We define $\mathcal{K}\coloneqq\{1,2,3\}$ and profile $q$ by
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\xi_{q}\coloneqq\left(\xi_{q,k})\right)_{i=1,2,3}}=\left(\begin{array}[]{c}3\\ 2\\ 1\\ \end{array}\right).$
Furthermore, we assume that we have an unlimited workforce available for each skill level.
Table 2: Set of possible skill compositions for profile $q$
$s_{1}$ $s_{2}$ $s_{3}$ $s_{4}$ $s_{5}$ $s^{i}_{q,1}$ 1 0 0 1 0 $s^{i}_{q,2}$ 1 2 1 0 0 $s^{i}_{q,3}$ 1 1 2 2 3
Table 2 contains the set $\mathcal{S}_{q}=\left\{s_{1},s_{2},s_{3},s_{4},s_{5}\right\}$ of skill compositions that can be used to assemble a team with profile $q$. We note that each skill composition $s_{i}$ is uniquely defined by their worker usage $\left(s^{i}_{q,k}\right)_{k=1,2,3}$ per skill level. For instance, the third column in Table 2 describes skill composition $s_{2}$, which uses two workers of skill level 2 and one worker of skill level 3 to assemble a team with profile $q$. In total, team route $(r,q)$ can be used to derive five unique team routes $\left\{(r,q,s_{i}):\ i=,1\dots,5\right\}$ in $\mathcal{R}^{\mathrm{D}}$.
We use these concepts to develop a model analogous to the AMP. For each $(r,q,s)\in\mathcal{R}^{\mathrm{D}}$, we define by $\beta_{k,\tau}^{r,s}(\omega_{\mathrm{max}})$ the number of workers of skill level $k$ required by route $r$ at time $\tau$, assuming worst-case travel times. These parameters take values $\beta_{k,\tau}^{r,s}=s_{q,k}$ for $\tau\in[\mathrm{tl}^{r,q},\mathrm{tr}^{r,q}]$ and $\beta_{k,\tau}^{r,s}=0$ otherwise. Using the previous definitions, we introduce the disaggregated master problem, abbreviated as DMP:
$\min$ $\sum_{(r,q,s)\in\mathcal{R}^{\mathrm{D}}}\mathbb{E}(c^{r})\lambda^{r}_{q,s}$ (10)
s.t. $\underset{(r,q,s)\in\mathcal{R}^{\mathrm{D}}:\ i\in\mathcal{I}^{r}}{\sum}\lambda^{r}_{q,s}\geq 1\quad\forall i\in\mathcal{I}$ (11)
$\underset{(r,q,s)\in\mathcal{R}^{\mathrm{D}}}{\sum}\beta_{k,\tau}^{r,s}(\omega_{\mathrm{max}})\lambda^{r}_{q,s}\leq N_{k}^{\mathrm{D}}\quad\forall k\in\mathcal{K},\ \forall\tau\in\mathcal{T}$ (12)
$\lambda^{r}_{q,s}\in\mathbb{N}\quad\forall(r,q,s)\in\mathcal{R}^{\mathrm{D}}$ (13)
Unlike in Section 4.1, constraints (12) consider workforces on an individual rather than an aggregated level. Similar to the ARMP, we call the DMP when only a subset of routes is considered and (13) is replaced by $\lambda^{r}_{q,s}\geq 0$ the aggregated reduced master problem (ARMP). For ease of reading, we will be referring to the columns of the DRMP as the set $\mathcal{R}^{\mathrm{D}}$. Moreover, we denote a solution $(\hat{\lambda}^{r}_{q})_{(r,q)\in\mathcal{R}}$ to the ARMP, which can not be extended to a solution $(\hat{\lambda}^{r}_{q,s})_{(r,q,s)\in{\mathcal{R}^{\mathrm{D}}}}$ of the DRMP as a disaggregated-infeasible solution. We note that this can be checked by solving the feasibility check, which has been proposed by Dall’Olio and Kolisch (2023) and is included in Appendix 10 of this publication. Analogous to Section 4.1, one can prove that there exists a binary solution for every feasible DMP instance.
Each disaggregated team route $(r,q,s)\in\mathcal{R}^{\mathrm{D}}$ corresponds to one column of the DRMP. It is easy to see that $|\mathcal{R}^{\mathrm{D}}|$ is considerably larger than $|\mathcal{R}|$. A formal proof of this is provided in Appendix 13. Additionally, each column in $\mathcal{R}$ can be obtained by convex combinations of columns in $\mathcal{R}^{\mathrm{D}}$. Thus, vertices of the feasible region of the linear relaxation of the ARMP may correspond to higher-dimensional faces in the DRMP. This makes it significantly harder to cut off non-integer solutions from the feasible region; therefore, solving the DRMP using a branch-and-cut algorithm becomes substantially harder. Experimental studies have shown that solving the DRMP is, on average, around 30% slower than solving the ARMP. Therefore, we only resort to solving the DRMP once a disaggregated-infeasible solution has been identified.
5 Solution Approach
In this section, we present a Branch-Price-Cut-and-Switch algorithm to solve the problem at hand. We initially start the algorithm at the root node with a subset of columns, denoted by $\bar{\mathcal{R}}\subset\mathcal{R}$, consisting of single-task tours $(d,i,d)$ for each task $i\in\mathcal{I}$, each profile $q\in\mathcal{Q}_{i}$ and the earliest possible depot leave time $tl\coloneqq\mathrm{ES}_{i}-t_{d,i}(\omega_{\mathrm{max}})$. To ensure feasibility, we include a column that finishes each task at the latest possible time $\mathrm{LF}_{i}^{\mathrm{e}}$ and uses up the entire available workforce. We then search the branching tree using a Branch-Price-and-Cut approach. If we obtain an integer solution, we perform the feasibility check proposed by Dall’Olio and Kolisch (2023) to check if the solution is disaggregated-feasible. If the solution fails the feasibility check, we mark the current node and its sibling node as disaggregated-infeasible and restart the procedure. We note that children of marked nodes are also marked by default. Whenever a disaggregated-infeasible node is encountered with the search tree, we solve the DRMP instead of the ARMP. The selection of unexplored nodes is done via a best-first search. Table 3 extends the notation summarized in Table 1 by model- and problem-specific notation introduced throughout Sections 4 and 5.
Table 3: Model- and problem-specific notation
Sets and Variables
$\hat{\mathcal{I}}_{q}$ Set of tasks that can be executed by profile $q$ including covering
$\mathcal{R}$ Set of feasible team routes for the AMP
$\bar{\mathcal{R}}$ Subset of all feasible team routes $\mathcal{R}$
$\mathcal{R}^{\mathrm{D}}$ Set of feasible team routes for the DMP
$\mathcal{G}$ Set of indices of rank-1 CGCs present at given node
$\lambda^{r}_{q}$ Binary variable, equals to 1 if feasible team route $(r,q)$ belongs to a solution of the AMP
$\lambda^{r}_{q,s}$ Binary variable, equals to 1 if feasible team route $(r,q,s)$ belongs to a solution of the DMP
Input and Model Parameters
$t_{i,j}^{\alpha}$ largest $\alpha$-quantile of the distribution of $t_{i,j}$
$b_{k,t}^{r,q}(\omega_{\mathrm{max}})$ Number of workers with at least level $k$ required by team route $(r,q)$ at time $t$ if $\omega_{\mathrm{max}}$ occurs
$\beta_{k,t}^{r,s}$ Number of workers with skill level $k$ required by team route $(r,q,s)$ at time $t$ if $\omega_{\mathrm{max}}$ occurs
$\mu_{i}$ Value of the dual variable associated with task covering constraints
$\delta_{k,t}$ Value of the dual variable associated with workforce constraints
$\psi_{g}$ Value of the dual variable associated with the rank-1 CGC with index $g$
$\mathcal{G}^{q}$ Pricing network in the ARMP for profile $q$
$\mathcal{G}^{q,s}$ Pricing network in the DRMP for disaggregated profile $(q,s)$
$\mathcal{A}^{q}$ Arcs of networks $\mathcal{G}^{q}$ and $\mathcal{G}^{q,s}$
$\mathcal{V}^{q}$ Nodes of networks $\mathcal{G}^{q}$ and $\mathcal{G}^{q,s}$
$W_{i,j}(F_{i}^{P,\mathrm{max}})$ time-dynamic weight of arc $(i,j)$ depending on the worst-case finish time $F_{i}^{P,\mathrm{max}}$ of the previous task
$F_{i}^{L,\mathrm{max}}$ worst-case finish time of task $i$ in the path associated with label $L$
$u_{i}$ rank-1 CGC coefficient for task constraint associated with task $i$
$u_{k,t}$ rank-1 CGC coefficient for workforce constraint associated with skill level $k$ and time $t$
In Section 5.1, we describe the pricing problem for the ARMP. Section 5.2 focuses on the labeling algorithm used to solve the pricing problem. In Section 5.3, necessary adjustments to solve the pricing problem associated with the DRMP are explained. Section 5.4 elaborates on several acceleration strategies used to speed up the solution process. Unless otherwise stated, all considerations can be transferred to the DRMP and its set $\mathcal{R}^{\mathrm{D}}$ of disaggregated team routes.
5.1 Pricing Problem for the ARMP
Whenever a solution to a reduced master problem (ARMP or DRMP) is obtained, we solve the corresponding pricing problem to check if there is a column with negative reduced costs. In the ARMP, the reduced cost of a column corresponding to a team route $(r,q)\in\mathcal{R}$ is equal to
$\mathbb{E}(c^{r})-\underset{i\in\mathcal{I}^{r}}{\sum}\mu_{i}-\underset{k\in\mathcal{K}}{\sum}\ \underset{\tau=\mathrm{tl}^{r,q}}{\overset{\mathrm{tr}^{r,q}}{\sum}}\delta_{k,\tau}b_{k,\tau}^{r,q}(\omega_{\mathrm{max}})$
where $\mu=(\mu_{i})_{i\in\mathcal{I}}$ and $\delta=(\delta_{k,\tau})_{k\in\mathcal{K},\tau\in\mathcal{T}}$ are the dual variables corresponding to constraints (11) and (12), respectively. We note that $\mu\geq 0$ and $\delta\leq 0$ holds.
The pricing problem of the ARMP corresponds to a set of elementary shortest path problems with resource constraints (ESPPRCs), one for each profile $q\in\mathcal{Q}$. These problems are typically solved using pricing networks, i.e., graphs where each path represents a column and its length is equal to the respective column’s reduced cost. We use a non-time expanded graph with time-dynamic arc weights, i.e., weights whose value depends on the current finish time distribution, to model and solve the associated pricing problems. We note that time-expanded graphs can also be used in this context, however they scale rather poorly for more granular time grids. Thus, they are less suited for the stochastic problem formulation at hand.
We now turn our focus to the construction of the graph used to solve the ESPPRC. First, we reiterate the concept of covering profiles introduced in Dall’Olio and Kolisch (2023). For profiles $q,\bar{q}\in\mathcal{Q}$ and a set $\mathcal{K}\coloneqq\{1,\dots,K\}$ of skill levels with $K\in\mathbb{N}_{>0}$, we say that $q$ covers $\bar{q}$ if
$\xi_{\bar{q},k^{*}}$ $=\xi_{q,k^{*}}\quad\text{and}$
$\xi_{\bar{q},k}$ $\leq\xi_{q,k}\quad\forall k\in\mathcal{K}$
hold, where $k^{*}=1$ is the skill level with the lowest qualifications. For a task $i\in\mathcal{I}$, we denote by $\hat{\mathcal{Q}_{i}}$ the set of profiles covering at least one profile in $\mathcal{Q}_{i}$. It might be advantageous to execute a task $i$ using a profile $q\in\hat{\mathcal{Q}_{i}}\setminus\mathcal{Q}_{i}$, e.g., in case of a shortage of low-skilled workers. Therefore, we denote by $\hat{\mathcal{I}}_{q}\coloneqq\{i\in\mathcal{I}:q\in\hat{\mathcal{Q}_{i}}\}$ the set of tasks that might be executed using profile $q$.
[…]
$\mathrm{ES}_{i}+p_{i,q}+t_{i,j}^{\alpha}>\mathrm{LF}_{j}-p_{j,q}$
holds, where
$t_{i,j}^{\alpha}\coloneqq\min\left\{t^{*}:\ \mathbb{P}(t_{i,j}\leq t^{*})\geq\alpha\right\}$
is the largest $\alpha$-quantile of the distribution of $t_{i,j}$. Therefore, we define the arc set $\mathcal{A}^{q}$ by
[…]
$\land\ \mathrm{ES}_{i}+p_{i,q}+t_{i,j}(\omega_{\max})\leq\mathrm{LF}_{j}^{\mathrm{e}}-p_{j,q}\}$
We can further reduce the size of $\mathcal{A}^{q}$ by removing arcs $(i,j)$ for which the following holds:
$\mathrm{ES}_{j}-\mathrm{LF}_{i}^{\mathrm{e}}$ $\geq d_{i,d}{(\omega_{\mathrm{max}})}+d_{d,j}{(\omega_{\mathrm{max}})}\quad$ (14)
If inequality (14) holds, replacing arc $(i,j)$ in a feasible route $r$ with arcs $(i,d)$ and $(d,j)$ splits $r$ into two new routes, which are also feasible and have the same joint objective function value as $r$. Furthermore, they do not occupy more workforce than $r$ at any time $t\in\mathcal{T}$. Hence, we can remove arc $(i,j)$ from $\mathcal{A}^{q}$.
Let $P$ be a path in $\mathcal{G}^{q}$ with depot leave time $\mathrm{tl}^{P}\geq 0$. Because the reduced costs of an arc $(i,j)\in\mathcal{A}^{q}$ are time-dependent and $\mathcal{G}^{q}$ does not contain any temporal information, arc weights are time-dynamic and specifically depend on the distribution of finish times $F_{i}^{P}$ and the worst-case finish time $F_{i}^{P}(\omega_{\mathrm{max}})$ at the previous node $i$ along $P$. For this, we set $w_{o^{\prime}}\coloneqq 0$, $\mu_{o^{\prime}}\coloneqq 0$ and $F_{o}^{P}(\omega_{\mathrm{max}})\coloneqq\mathrm{tl}^{P}-1$. Then, the weight $W_{i,j}\left(F_{i}^{P}(\omega_{\mathrm{max}})\right)$ of an arc $(i,j)\in\mathcal{A}^{q}$ is given by
$W_{i,j}\left(F_{i}^{P}(\omega_{\mathrm{max}})\right)\coloneqq w_{j}\cdot\mathbb{E}(F_{j}+P_{j}(F_{j}))-\mu_{j}-\underset{k\in\mathcal{K}}{\sum}\ \underset{\tau=F_{i}^{P}(\omega_{\mathrm{max}})+1}{\overset{F_{j}^{P}(\omega_{\mathrm{max}})}{\sum}}\delta_{k,\tau}b_{k,\tau}^{P,q}(\omega_{\mathrm{max}}).$
Then, finding a minimum-cost $o-o^{\prime}$-path that satisfies constraints (3) and (4) is equivalent to finding a feasible team route $(r,q)\in\mathcal{R}$ (i.e., a new column) with minimum reduced cost.
5.2 Labeling Algorithm for the ARMP Pricing Problem
[…]
$\mathrm{tl}^{L^{\prime}}$ $=\mathrm{tl}^{L}$
$P^{\prime}$ $=(o,v_{1},\dots,v_{l},v,v^{\prime})$
$T_{v^{\prime}}^{\mathrm{cost}}$ $=T_{v}^{\mathrm{cost}}+W_{v,v^{\prime}}(F_{v}^{L,\mathrm{max}})$
$T_{v^{\prime},i}^{\mathrm{perf}}$ $=\begin{cases}\begin{aligned} &T_{v,i}^{\mathrm{perf}}-1&&\text{if}\ \ i=v^{\prime}\\ &1&&\text{if}\ \ F_{v^{\prime}}^{L^{\prime},\mathrm{max}}+t_{v^{\prime},i}(\omega_{\mathrm{max}})>\mathrm{LF}_{i}^{\mathrm{e}}-p_{i,q}\\ &1&&\text{if}\ \ F_{v^{\prime}}^{L^{\prime},\mathrm{max}}+t_{v^{\prime},i}^{\alpha}>\mathrm{LF}_{i}-p_{i,q}\\ &T_{v,i}^{\mathrm{perf}}&&\text{otherwise}\end{aligned}\end{cases}$
$F_{v^{\prime}}^{L^{\prime}}(\tau)$ $=\begin{cases}\begin{aligned} &\underset{z=0}{\overset{\mathrm{ES}_{v^{\prime}}+p_{v^{\prime},q}}{\sum}}\mathbb{P}(F_{v}^{L}+t_{v,v^{\prime}}+p_{v^{\prime},q}=z)&&\text{if}\ \ \tau=\mathrm{ES}_{v^{\prime}}+p_{v^{\prime},q}\\ &\mathbb{P}(F_{v}^{L}+t_{v,v^{\prime}}+p_{v^{\prime},q}=\tau)&&\text{if}\ \ \tau>\mathrm{ES}_{v^{\prime}}+p_{v^{\prime},q}\\ &0&&\text{otherwise}\end{aligned}\end{cases}$
$F_{v^{\prime}}^{L^{\prime},\mathrm{max}}$ $=\max\left\{F_{v}^{L,\mathrm{max}}+t_{v,v^{\prime}}(\omega_{\mathrm{max}}),\mathrm{ES}_{v^{\prime}}\right\}+p_{v^{\prime},q}$
where we define $p_{o^{\prime},q}=0$. We note that we reset the resources $T_{v^{\prime},i}^{\mathrm{perf}}$ of tasks that cannot be visited anymore without violating time windows or chance constraints to 1, as this strengthens the dominance relations between labels. Furthermore, $\mathrm{tl}^{L},v$ and $P$ are properties necessary for a label $L$ to be well-defined, however they are not resources in the classical sense.
We say that the extension of label $L$ along $(v,v^{\prime})$ is feasible if the resulting label $L^{\prime}$, which is obtained from $L$ using the above resource extension function, is feasible for $\mathcal{G}^{q}$.
A core component of every labeling algorithm is its dominance rule, which allows the discarding of labels that can not be part of an optimal $oo^{\prime}$ path.
Definition 5.1 (Dominance Rule)
[…]
$T_{v}^{1,\mathrm{cost}}$ $\leq T_{v}^{2,\mathrm{cost}}$ (15)
[…]
$\mathbb{P}(F_{v}^{1}\leq\tau)$ $\geq\mathbb{P}(F_{v}^{2}\leq\tau)\quad\forall\tau\in[\mathrm{ES}_{v}+p_{v,q},\mathrm{LF}_{v}^{\mathrm{e}}]$ (17)
$F_{v}^{1,\mathrm{max}}$ $=F_{v}^{2,\mathrm{max}}$ (18)
Properties (16) and (17) ensure that any feasible extension of $L^{2}$ is also feasible for $L^{1}$. Constraints (15), (17) and (18) guarantee that, after extending both $L^{1}$ and $L^{2}$ along $(v,v^{\prime})$, the reduced cost of $L^{1^{\prime}}$ is still less or equal to the reduced cost of $L^{2^{\prime}}$.
[…]
$\left[\mathrm{ES}_{i}-t_{o,i}(\omega_{\mathrm{max}}),\mathrm{LF}_{i}-p_{i,q}-t_{o,i}^{\alpha}\right].$ (19)
Leaving the depot at a time instant greater than $\mathrm{LF}_{i}-p_{i,q}-t_{o,i}^{\alpha}$ would violate chance constraint (3) at task $i$, while leaving before $\mathrm{ES}_{i}-t_{o,i}(\omega_{\mathrm{max}})$ incurs unnecessary waiting time. These labels are also called initial labels. We then extend these labels using the previously described resources extension function, discard dominated labels, and repeat the same procedure for the oldest label until no feasible extensions can be found anymore. We note that the calculation of the start time distribution consumes the majority of runtime during label extensions. When the support of travel time distributions is small enough, these calculations can be done exactly. We refer to Errico et al. (2016) and Errico et al. (2018) for an extensive description of such an algorithm.
5.3 Peculiarities for the DRMP
Though the DRMP is structurally very similar to the ARMP, several characteristics of the former must be considered when solving the pricing problem of the DRMP. Because the node and arc set does not depend on the skill composition $s$, we can define the graph of a disaggregated profile $(q,s)\in\mathcal{Q}^{\mathrm{D}}$ as $\mathcal{G}^{q,s}\coloneqq(\mathcal{V}^{q},\mathcal{A}^{q},\beta)$, where the only difference lies in the coefficients $\beta_{k,t}^{s}(\omega_{\mathrm{max}})$ replacing $b_{k,t}^{q}(\omega_{\mathrm{max}})$ during the calculation of the time-dynamic arc weights. We note that $\beta\coloneqq\left(\beta_{k,t}^{s}(\omega_{\mathrm{max}})\right)_{k\in\mathcal{K},t\in\mathcal{T}}$ denotes the disaggregated workforce requirements of a team with profile $q$ and skill composition $s$.
A significant difference lies in the number of pricing networks. While there is exactly one pricing network for each profile $q\in\mathcal{Q}$ for the ARMP, there is one pricing network for each disaggregated profile $(q,s)\in\mathcal{Q}^{\mathrm{D}}$. Typically, the cardinality of $\mathcal{Q}^{\mathrm{D}}$ grows exponentially in $|\mathcal{Q}|$. Thus, creating and solving a unique pricing network for each disaggregated profile could render a column generation approach highly impractical due to the vast number of networks to be solved. In the following, we show that multiple pricing networks can be solved simultaneously when the used dominance rule is slightly altered.
Let $(q,s)\in\mathcal{Q}^{\mathrm{D}}$ be a disaggregated profile. We then introduce an adjusted dominance rule:
Definition 5.2 (Dominance rule for the DRMP)
Let
[…]
be labels for $k=1,2$. We say that $L^{1}$ dominates $L^{2}$ in $\mathcal{G}^{q,s}$ if the following properties hold:
$T_{v}^{1,\mathrm{cost}}+\underset{k\in\mathcal{K}}{\sum}\underset{\tau=\mathrm{tl}^{1}}{\overset{F_{v}^{1,\mathrm{max}}}{\sum}}\delta_{k,\tau}\beta_{k,\tau}^{P^{1},s,1}(\omega_{\mathrm{max}})$ $\leq T_{v}^{2,\mathrm{cost}}+\underset{k\in\mathcal{K}}{\sum}\underset{\tau=\mathrm{tl}^{2}}{\overset{F_{v}^{2,\mathrm{max}}}{\sum}}\delta_{k,\tau}\beta_{k,\tau}^{P^{2},s,2}(\omega_{\mathrm{max}})$ (20)
[…]
$\mathbb{P}(F_{v}^{1}\leq\tau)$ $\geq\mathbb{P}(F_{v}^{2}\leq\tau)\quad\forall\tau\in[\mathrm{ES}_{v}+p_{v,q},\mathrm{LF}_{v}^{\mathrm{e}}]$ (22)
$F_{v}^{1,\mathrm{max}}$ $=F_{v}^{2,\mathrm{max}}$ (23)
$\mathrm{tl}^{1}$ $\geq\mathrm{tl}^{2}$ (24)
Dominance rule 5.2 offsets the reduced costs of $L^{1}$ and $L^{2}$ by the workforce penalty of each path and imposes an additional restriction (24) on their depot leave times. It is easy to see that if $L^{1}$ dominates $L^{2}$ in $\mathcal{G}^{q,s}$, it also dominates $L^{2}$ with respect to the dominance rule described in Definition 5.1. However, the converse is not true.
Figure 2: Labels satisfying dominance rule 5.1 and violating rule 5.2
Example 5.3
Figure 2 visualizes the pricing network for a disaggregated profile $(q,s)\in\mathcal{Q}^{\mathrm{D}}$ and two labels $L^{1}$ (solid line) and $L^{2}$ (dotted line). We assume $\mathcal{K}=\{1\}$, $s_{q,1}=1$ and $\delta_{1,\tau}=-1$ for all $\tau\in[0,10]$. Furthermore, we simplify our considerations by neglecting time window restrictions and setting $w_{1}=w_{2}=1$. Moreover, we define $p_{2,q}=3$ and $p_{1,q}=3$. The reduced costs for labels $L^{1}$ and $L^{2}$ are then equal to $T_{1}^{1,\mathrm{cost}}=13.5$ and $T_{2}^{2,\mathrm{cost}}=18$. Therefore, it is clear that $L^{1}$ dominates $L^{2}$ in the sense of Definition 5.1. However, if we offset the reduced cost of each label by the workforce penalties, we obtain
$T_{1}^{2,\mathrm{cost}}+\underset{\tau=0}{\overset{10}{\sum}}\delta_{1,\tau}\beta_{k,\tau}^{P^{2},s,2}(\omega_{\mathrm{max}})=18-11=7<7.5=13.5-6=T_{1}^{1,\mathrm{cost}}+\underset{\tau=5}{\overset{10}{\sum}}\delta_{1,\tau}\beta_{k,\tau}^{P^{1},s,1}(\omega_{\mathrm{max}})$
Hence, $L^{1}$ does not dominate $L^{2}$ in $\mathcal{G}^{q,s}$, i.e., in the sense of Definition 5.2. The dominance rule presented here tends to be more strict, leading to around 25% less dominated labels.
[…]
that is, all properties besides the reduced cost $T_{v}^{\mathrm{cost}}$ of $L$ remain the same. As the dominance rule on $\mathcal{G}^{q,s}$ merely considers reduced costs after offsetting them by the workforce penalty, it is clear that if a label $L^{1}$ dominates another label $L^{2}$ in $\mathcal{G}^{q,s}$ if and only if it dominates $L^{2}$ in $\mathcal{G}^{q,\tilde{s}}$ for all $\tilde{s}\in\mathcal{S}_{q}$. For a formal proof of these relations, see Appendix 12.
Therefore, for a profile $q\in\mathcal{Q}$, the column with minimum reduced cost for the set $\left((q,s)\right)_{s\in\mathcal{S}_{q}}$ of disaggregated profiles can be calculated as follows: we select an arbitrary skill composition $\tilde{s}\in\mathcal{S}_{q}$ and solve the ESPPRC on $\mathcal{G}^{q,\tilde{s}}$. After no further labels can be created, the following optimization problem on the set $\mathcal{L}$ of non-dominated labels present at the sink $o^{\prime}$ is solved to obtain a route with profile $q$, an optimal skill composition $s^{*}$ and minimal reduced cost:
$\min\left\{T^{L,\mathrm{cost}}_{o^{\prime}}+\underset{\tau=\mathrm{tl}^{L}}{\overset{\mathrm{tr}^{L}}{\sum}}\delta_{k,\tau}\beta_{k,\tau}^{L,\tilde{s}}-\underset{\tau=\mathrm{tl}^{L}}{\overset{\mathrm{tr}^{L}}{\sum}}\delta_{k,\tau}\beta_{k,\tau}^{L,s}:\ s\in\mathcal{S}_{q},\ L\in\mathcal{L}\right\}$
Because the number of labels at the sink and the number of possible skill compositions is usually small, this problem can be solved by explicit enumeration. In total, the runtime savings by solving one pricing network (instead of multiple) for each profile exceeds the additional runtime caused by a weaker dominance rule by several orders of magnitude.
We note that the above approach is also applicable to general SPPRCs if several properties are satisfied. Let $\mathcal{N}$ be a set of pricing networks and $\mathcal{N}=\underset{i=1}{\overset{n}{\bigcap}}\mathcal{N}_{i}$ with $n\in\mathbb{N}$ be a disjunct partition of $\mathcal{N}$. Then, one can apply the above methodology to a subset $\mathcal{N}_{i}$ of pricing networks for an arbitrary $i\in\mathcal{N}$ if the following properties are satisfied:
  • i)
    a path $P$ induced by a label $L$ in a pricing network $\mathcal{G}$ in $\mathcal{N}_{i}$ is feasible in $\mathcal{G}$ if and only if it is feasible in all pricing networks in $\mathcal{N}_{i}$,
  • ii)
    there exists an adjusted dominance rule that only depends on properties shared among all pricing networks in $\mathcal{N}_{i}$, i.e., it is agnostic of the specific pricing network in $\mathcal{N}_{i}$ that is considered, and
  • iii)
    if a label $L$ dominates a label $\tilde{L}$ with respect to the adjusted dominance rule, one can prove that it dominates $\tilde{L}$ in all pricing networks in $\mathcal{N}_{i}$ with respect to the original dominance rule.
In the application at hand, the set of pricing networks is given by $\mathcal{N}\coloneqq\left\{\mathcal{G}^{q,s}:\ q\in\mathcal{Q},s\in\mathcal{S}^{q}\right\}$ and its partition is $\mathcal{N}_{q}\coloneqq\left\{\mathcal{G}^{q,s}:\ s\in\mathcal{S}^{q}\right\}$ for all $q\in\mathcal{Q}$. Furthermore, path feasibility is defined by the satisfaction of constraints (3)–(4), which do not depend on the skill composition, i.e., (i) is satisfied by the definition of $\mathcal{N}_{q}$. Moreover, it is clear to see that the adjusted dominance rule (20)–(24) does not depend on the skill composition, which is the only property that is not shared among all pricing networks within $\mathcal{N}_{q}$. Finally, the usage of offset reduced costs (20), jointly with inequality (24), allows for the proof of dominance with respect to the original dominance rule (15)–(18) in all pricing networks in $\mathcal{N}_{q}$ if dominance with respect to (20)–(24) is satisfied in an arbitrary pricing network in $\mathcal{N}_{q}$. For additional details on how the equivalence of pricing networks can be derived using properties i)–iii), we refer to the proof presented in Appendix 12.
5.4 Acceleration Strategies
As the pricing step is by far the most runtime-intensive component in the proposed solution approach, we employ several strategies that allow us to generate columns with negative reduced costs more quickly. In the following, we describe these techniques in detail.
Graph Size Reduction
We use two heuristics that initially reduce the graph size and set of initial labels and gradually increase them when necessary. The first heuristic consists of separating the set of possible depot leave times (19) into bins {1,…,B} of equal size and sorting these bins in an ascending order with respect to the minimum reduced cost of all labels inside the bin. We then solve the ESPPRC considering only the initial labels in bin $1$. If no column with negative reduced costs has been found, we continue with the next bin until we either find a negative column or solve the final bin $B$. In our algorithm, we set the value $B$ such that each bin $\{1,\dots,B-1\}$ has a size of $5$.
[…]
In our experiments, we set $\Delta=4$ as this returned the best results. Furthermore, we first iterate through all initial label bins and then switch to allowing extensions along all arcs only if no negative column has been found in any of the bins.
Decremental State Space Relaxation
[…]
Moreover, Righini and Salani (2008) and Boland, Dethridge, and Dumitrescu (2006) present several other methods to solve the ESPPRC, which rely on relaxing task resources and only adding them if needed. Righini and Salani (2008) develop a branch-and-bound framework, while Boland, Dethridge, and Dumitrescu (2006) propose several algorithms to forbid node repetitions. We note that in the test instances studied, time windows are often so tight that we rarely observed cycles, hence implementing such advanced methods to deal with non-elementary paths does not have a significant impact on the algorithm’s performance for the problem at hand.
5.5 Branching Strategies
We use three different branching rules to cut off fractional solutions of the ARMP or DRMP. In the following, we summarize these strategies and their technicalities.
Branching on Task Finish Times
Gélinas et al. (1995) were among the first ones to branch on resource windows, more specifically, time windows. While this strategy is not novel, the process of selecting a task and a time instant to branch on has a significant impact on the branching rule’s efficiency. Let $(\bar{\lambda}^{r})_{r\in\mathcal{R}}$ be an optimal solution to the ARMP and let $\mathcal{R}_{\mathrm{F}}\subseteq\mathcal{R}$ be the set of columns for which $\bar{\lambda}^{r}$ is fractional. For each task $i\in\mathcal{I}$, we define
$\mathcal{B}^{\mathrm{F}}(i)\coloneqq\left\{F_{i}^{r,q}(\omega_{\mathrm{max}}):\ (r,q)\in\mathcal{R}_{\mathrm{F}}:\ i\in r\right\},$
$\mathcal{B}^{\mathrm{F}}_{U}(i)\coloneqq\left\{F_{i}^{r,q}(\omega_{\mathrm{max}}):\ (r,q)\in\mathcal{R}_{\mathrm{F}}:\ i\in r,\ F_{i}^{r,q}(\omega_{\mathrm{max}})\ \text{unique}\right\}$
as the set of all (unique) worst-case finish times of task $i$ within team routes with fractional values, respectively. We then select all tasks for which the cardinality of the latter set is maximal and denote it by $\mathcal{I}^{\mathrm{F}}$, i.e.,
$\mathcal{I}^{\mathrm{F}}=\left\{i\in\mathcal{I}:\left|\mathcal{B}^{\mathrm{F}}_{U}(i)\right|=\max\left\{\ \left|\mathcal{B}^{\mathrm{F}}_{U}(j)\right|:\ j\in\mathcal{I}\right\}\right\}.$
We then select the task $i^{*}$ for which the standard deviation of the worst-case finish times of task $i^{*}$ in all routes in $\mathcal{B}^{\mathrm{F}}(i)$ is maximal, that is
[…]
where $\sigma(\cdot)$ is the standard deviation of a finite set. We then branch on task $i^{*}$ and time instant $\tau^{*}=\left\lfloor M(\mathcal{B}^{\mathrm{F}}(i^{*}))\right\rfloor$, where $M(\cdot)$ is the median of a finite set. We then create two child nodes and impose constraints
$F_{i^{*}}^{r,q}(\omega_{\mathrm{max}})\leq\tau^{*}\quad\forall(r,q)\in\mathcal{R}:\ i^{*}\in r\quad\text{or}\quad F_{i^{*}}^{r,q}(\omega_{\mathrm{max}})>\tau^{*}\quad\forall(r,q)\in\mathcal{R}:\ i^{*}\in r,$
respectively. Furthermore, we remove all tours that violate the constraints from the child nodes. In the pricing steps, we forbid extensions that would violate these task finish time constraints. If
$\max\left\{\sigma\left(\left\{F_{i}^{r,q}(\omega_{\mathrm{max}}):\ (r,q)\in\mathcal{B}^{\mathrm{F}}(i)\right\}\right):\ i\in\mathcal{I}^{\mathrm{F}}\right\}=0$
holds, we know by construction that in the current solution, $\left|\mathcal{B}^{\mathrm{F}}_{U}(i)\right|=1$ holds for all $i\in\mathcal{I}$. Therefore, independent of the selection of $i^{*}$ and $\tau^{*}$, the above branching rule produces a trivial branch. If this is the case, a different branching strategy must be used. We note that the above branching strategy neither changes the master problems nor the pricing problem’s structure.
Branching on Number of Tours at a Given Time
Desrochers, Desrosiers, and Solomon (1992) introduced branching on tour counts, which has proven to be efficient for vehicle routing problems. Let $\tau^{*}$ be the time instant for which $l_{\tau^{*}}=\underset{(r,q)\in\mathcal{R}}{\sum}g_{\tau^{*}}^{r}\lambda^{r}_{q}$ is closest to 0.5, where $g_{\tau}^{r}=1$ if $\mathrm{tl}^{r,q}\leq g_{\tau}^{r}\leq\mathrm{tr}^{r,q}$ and $g_{\tau}^{r}=0$ else. We then branch on the current solution by imposing
$\underset{(r,q)\in\mathcal{R}}{\sum}g_{\tau^{*}}^{r}\lambda^{r}_{q}\leq\lfloor l_{\tau^{*}}\rfloor\quad\text{or}\quad\underset{(r,q)\in\mathcal{R}}{\sum}g_{\tau^{*}}^{r}\lambda^{r}_{q}\geq\lceil l_{\tau^{*}}\rceil.$ (25)
The pricing problem has to be adjusted in the sense that time-dynamic arc weights have to consider the dual cost of the additional inequalities. If a label $L$ is extended along an arc $(v,v^{\prime})$ and $F_{v^{\prime}}^{L}(\omega_{\mathrm{max}})<\tau^{*}\leq F_{v^{\prime}}^{L}(\omega_{\mathrm{max}})$ holds, the cost of arc $(v,v^{\prime})$ must be adjusted by the dual cost of (25). Similar to branching on task finish times, this branching strategy can potentially produce a trivial branch, making a fallback branching option necessary.
Branching on Variables
If both previously described branching rules fail to produce a nontrivial branching decision, we select the most fractional variable $\lambda^{r^{*}}_{q^{*}}$ and set it to 0 or $1$ in the child nodes. When a team route $(r^{*},q^{*})$ is forced, i.e., $\lambda^{r^{*}}_{q^{*}}=1$ is enforced, we remove all tasks which are visited by $r^{*}$ from the pricing networks, discard all team routes that share tasks with $r^{*}$ and adequately reduce the total available workforce $N_{k}$ and $N_{k}^{\mathrm{D}}$ for all $k\in\mathcal{K}$. When a team route $(r^{*},q^{*})$ is forbidden, we introduce an additional resource to the pricing problem that tracks how many arcs a label has used that correspond to segments of $r^{*}$. If this resource is fully consumed, we discard the label as it equals a forbidden tour. For the DRMP, we do not discard said labels but skip the skill composition $s$ that is used to execute the forbidden route. Furthermore, if a label is equal to a subpath of a forbidden tour, it can not dominate any other label as this might lead to cutting off optimal labels.
5.6 Cutting Planes
Whenever a fractional optimal solution is found at a node in the branching tree, we model the problem of finding a most violated rank-1 Chvátal-Gomory cut (CGC) as a mixed-integer problem and solve it using a generic solver with a very short time limit. If optimality could not be proven within the time limit, the best solution found so far is returned. If no solution is found, the cutting plane step is skipped and the algorithm proceeds with a branching step. This approach was first proposed by Fischetti and Lodi (2007) and applied to a vehicle routing problem with time windows by Petersen, Pisinger, and Spoorendonk (2008). In the following, we base our considerations on the ARMP formulation. Let $(u_{i})_{i\in\mathcal{I}}\in[0,1)$ and $(u_{k,t})_{k\in\mathcal{K},t\in\mathcal{T}}$ be the coefficients of the most violated cut. We then add the constraint
$\underset{(r,q)\in\mathcal{R}}{\sum}\left\lfloor\underset{i\in\mathcal{I}^{r}}{\sum}u_{i}+\underset{k\in\mathcal{K}}{\sum}\ \underset{\tau=\mathrm{tl}^{r,q}}{\overset{\mathrm{tr}^{r,q}}{\sum}}b_{k,\tau}^{r,q}(\omega_{\mathrm{max}})u_{k,t}\right\rfloor\lambda^{r}_{q}\leq\left\lfloor\underset{i\in\mathcal{I}^{r}}{\sum}u_{i}+\underset{k\in\mathcal{K}}{\sum}N_{k}\cdot\left(\underset{\tau\in\mathcal{T}}{\sum}u_{k,t}\right)\right\rfloor$
to the master problem and re-solve the current node.
[…]
$T_{v^{\prime}}^{\mathrm{cost}}$ $=T_{v}^{\mathrm{cost}}+\left\lfloor T_{v,g}+u_{v^{\prime}}^{g}+\underset{k\in\mathcal{K}}{\sum}\ \underset{\tau=F_{v}^{L,\mathrm{max}}}{\overset{F_{v^{\prime}}^{L^{\prime},\mathrm{max}}}{\sum}}b_{k,\tau}^{P,q}(\omega_{\mathrm{max}})u_{k,t}^{g}\right\rfloor\cdot\psi_{g}$
$T_{v^{\prime},g}$ $=T_{v,g}+u_{v^{\prime}}^{g}+\underset{k\in\mathcal{K}}{\sum}\ \underset{\tau=F_{v}^{L,\mathrm{max}}}{\overset{F_{v^{\prime}}^{L^{\prime},\mathrm{max}}}{\sum}}b_{k,\tau}^{P,q}(\omega_{\mathrm{max}})u_{k,t}^{g}-\left\lfloor T_{v,g}+u_{v^{\prime}}^{g}+\underset{k\in\mathcal{K}}{\sum}\ \underset{\tau=F_{v}^{L,\mathrm{max}}}{\overset{F_{v^{\prime}}^{L^{\prime},\mathrm{max}}}{\sum}}b_{k,\tau}^{P,q}(\omega_{\mathrm{max}})u_{k,t}^{g}\right\rfloor$
where $\psi_{g}\geq 0$ is the dual variable associated with CGC $g\in\mathcal{G}$. Dominance rule 5.1 is then slightly adjusted and extended:
Definition 5.4 (Dominance Rule)
Let
[…]
be labels for $k=1,2$. Let $\mathcal{G}$ be the set of indices of CGCs and $\mathcal{G}^{>}\coloneqq\left\{g\in\mathcal{G}:\ T_{v,g}^{1}>T_{v,g}^{2}\right\}$. We say that $L^{1}$ dominates $L^{2}$ if constraints (16)–(18) and
$T_{v}^{1,\mathrm{cost}}+\underset{g\in\mathcal{G}^{>}}{\sum}\psi_{g}$ $\leq T_{v}^{2,\mathrm{cost}}$ (26)
$T_{v,g}^{1}$ $\leq T_{v,g}^{2}\quad\forall g\in\mathcal{G}\setminus\mathcal{G}^{>}$ (27)
hold.
Inequalities (26) and (27) ensure that for any feasible extension $(v,v^{\prime})$ of $L^{2}$, the reduced costs still remain dominated by $L^{1}$. For a proof of correctness of this approach, we refer to Section 4 of Petersen, Pisinger, and Spoorendonk (2008).
Identifying violated CGCs can be quite costly if repeated frequently. Furthermore, each CGC slightly weakens the dominance rule due to the additional constraints (27), and calculating the coefficients $T_{v,g}$ requires additional computational effort. Hence, limiting the maximum amount of CGC present at any node in the search tree can be beneficial. Finally, we note that the above CGCs remain feasible when the master problem formulation is switched to the DRMP.
5.7 An Early Termination Heuristic
Because several exterior factors, such as unexpected aircraft delays, can create the need for re-optimization, limiting the maximum runtime of the Branch-Price-Cut-and-Switch scheme is often necessary. Let $\bar{\mathcal{R}}$ and $\bar{\mathcal{R}^{\mathrm{D}}}$ be the sets of columns found during the branch-and-price procedure for the ARMP and DRMP, respectively. We then define the set $\mathcal{R}^{\mathrm{D},\mathrm{H}}$ as
$\mathcal{R}^{\mathrm{D},\mathrm{H}}\coloneqq\left\{(r,q,\tilde{s}):(r,q)\in\bar{\mathcal{R}},\tilde{s}\in\mathcal{S}_{q}\right\}\ \cup\ \left\{(r,q,\tilde{s}):\exists s\in\mathcal{S}_{q}:\ (r,q,s)\in\bar{\mathcal{R}^{\mathrm{D}}},\tilde{s}\in\mathcal{S}_{q}\right\}$
Whenever the time limit is reached, we solve the DMP (10)–(13), including integrality constraints, on the column set $\mathcal{R}^{\mathrm{D},\mathrm{H}}$. If a feasible solution has been found within a prespecified time limit, we return the solution as the best integer solution found.
6 Experimental Study
In the following, we analyze the impact of cutting planes, switching between the DRMP and ARMP as master problem formulations, and different branching strategies on solution quality and convergence speed. Furthermore, we compare stochastic and deterministic solutions and evaluate the impact of stochasticity on optimal strategies.
Section 6.1 elaborates on the generation of test instances. Section 6.2 summarizes the algorithm’s performance for different configurations of the aforementioned components. In Section 6.3, stochastic and deterministic optimal policies are compared with respect to their practical feasibility. The algorithm is implemented using Python 3.11 and Gurobi 12.0.0. Furthermore, all computational studies were performed on a single machine equipped with an Intel® Xeon® W-1390p 11th gen 8-core 3.5GHz processor, 32GB of RAM and running Windows 10. All instances and corresponding solutions can be found under https://github.com/andreashagntum/StochasticTeamFormationRoutingAirport.
6.1 Instance Set
We use the instance generator developed by Dall’Olio and Kolisch (2023) to construct a set of test instances of different complexity. A predefined number of flights and their characteristics, such as time windows and parking positions, are generated based on realistic assumptions from Munich Airport. In the following, we briefly elaborate on parameters that have been varied throughout the generation.
In order to generate instances of different sizes, we vary the length of the planning horizon between 60, 90, and 120 minutes and generate 10, 20, and 30 flights per hour. The underlying time grid consists of equidistant time steps with a length of 2 minutes. Each type of airplane served at Munich Airport can be loaded using one of up to 3 different team formations, called slow, intermediate, and fast mode. Faster modes use up more workers but require less time to (un)load an aircraft. Not every plane can be loaded with all 3 modes, especially smaller planes typically only support slow or fast modes. We note that the same properties hold for unloading tasks.
We generate multiple instance sets by restricting the available team formations to only the intermediate mode, to the slow and the fast mode, or to all three modes, respectively. Furthermore, the available workforce ranges between 10% and 90% of the workers required when all tasks are started at the earliest possible time with the fastest possible mode. In the following, this factor is referred to as worker strength. Lastly, the minimum service level is set to $\alpha=0.9$ and the extended latest finish time is set to $\mathrm{LF}_{i}^{\mathrm{e}}=\mathrm{LF}_{i}+5$, i.e., 5 time steps after the latest finish time. We generated five random flight schedules for each combination of the aforementioned parameters, leading to a total of $3\cdot 3\cdot 3\cdot 9\cdot 5=1,215$ instances. We note that the largest test instances in this set, namely instances with a 2-hour planning horizon, 30 flights per hour, and three available modes, replicate the most complex instances one might encounter at Munich Airport, which are only observed during on-peak times. Furthermore, at Munich Airport, workforce planning is done based on 8-hour shifts, each of which includes up to two on-peak slots of at most 2 hours in length. Hence, during off-peak times, there is sufficient workforce available to decide on team formations and task assignments in a straightforward fashion, i.e., each task is executed by a team, which immediately returns back to the depot after finishing said task. We note that this is the approach currently taken at Munich Airport. However, during on-peak times, one needs to ensure that the available workforce is utilized efficiently in order to execute all scheduled tasks within their specified time windows.
All further properties, such as flight schedules and task execution times, are calculated as described by Dall’Olio and Kolisch (2023). For a precise description of the process, we refer to Section 6.1 and Appendices E and F of their paper.
We approximate the distribution of travel times between locations using a quadratic regression model that aims to predict travel times based on several exogenous factors. For this purpose, we collected real-world data on travel times of various ground operating vehicles, such as passenger buses and baggage handling vehicles, at Terminal 2 of Munich Airport for the entire year of 2024. These values, if needed, are adjusted to baggage handling vehicles by scaling them proportionally by their respective average speeds. We then categorized all datapoints based on several features, which in previous studies were identified to have a significant impact on travel times: weather conditions (cloudy, sunny, rainy, snowy), time during the day (on-peak or off-peak) and the week (weekday or weekend, holiday season yes or no), number of required apron crossings to get from origin to destination gate, distance, and best-case travel time. To estimate travel time distributions for a given gate pair under specified external conditions, we then apply the quadratic regression model to predict the percentual deviation from the shortest travel time. To account for variability, we then identify historical observations from other gate pairs with similar deterministic travel times, distances, and numbers of crossings. For these similar observations, we calculate the prediction errors of the percentual delay under their respective external conditions. Finally, we construct the travel time distribution for the target gate pair as follows. For each residual obtained from the aforementioned similar observations, we add the residual to the predicted percentual delay for the target gate pair and multiply the result by the best-case travel time. The resulting set of datapoints is then used to derive an empirical distribution.
For the test instances used in this section, we assumed an intra-day on-peak time on a sunny weekday during holiday season. In practice, this setting resembles a typical high-demand day at the beginning of summer holiday season.
6.2 Influence of the Algorithm’s Features
From a methodological point of view, our solution approach differs from the ones typically used for variants of vehicle routing or team formation problems in three ways: dynamically switching between two master problem formulations, branching on task finish times and an exact separation of rank-1 CGCs. In the following, these are called the algorithm’s features. In this section, we compare a total of six solver configurations: we first enable and then disable all three features, resulting in what are called ‘full’ and ‘basic’ configurations (or algorithms), respectively. Furthermore, we implemented the branch-and-price approach proposed by Yuan, Liu, and Jiang (2015), which differs from our approach in several ways. First, they do not use cutting planes and employ branching on arcs as their sole branching rule. Second, they apply a greedy and a neighborhood search heuristic to find negative columns before applying a labeling algorithm. Third, they only use a disaggregated master problem formulation, i.e., they do not switch between master problem formulations, and always return all negative columns found rather than just the most negative column. Fourth, while we aggregate multiple pricing networks into one by comparing offset reduced costs and adding an additional restriction (24) on the depot leave times, they demand reduced cost dominance for all skill compositions. Thus, unlike us, they implicitly solve all pricing networks exactly.
Moreover, we disable exactly one feature to obtain three additional variants of our solution method called ‘no DRMP’, ‘no CGCs’ and ‘no branching on task finish times’ (also abbreviated as ‘no branching’). For instance, the configuration ‘no DRMP’ uses branching on task finish times and adds up to 12 CGCs at the root node, but does not dynamically switch to the DRMP master problem formulation. We note that we do not include solver configurations with exactly two features enabled in our analysis, as the results do not fundamentally differ from the ones presented in the following.
Feature Design
We configure the aforementioned features as follows. When branching, we first try to branch on task finish times by using the procedure described in Section 5.5. If that is not possible, we look for branches on vehicle counts and, if this also fails, we use the fallback option of branching on variables. Furthermore, whenever a disaggregated-infeasible solution is found, we switch to solving the DRMP formulation at the current node and its sibling node. Additionally, child nodes of DRMP nodes inherit their master problem’s type, i.e., they are also solved using the DRMP formulation. Finally, we only search for violated CGCs at the root node and add up to 12 cuts. The MIP required to identify such cuts, as described by Fischetti and Lodi (2007) is solved using Gurobi with a time limit of 0.3 seconds. For each instance, a hard time limit of 180 seconds is imposed. If no optimal solution has been found, an upper bound is obtained by the procedure described in Section 5.7. We note that preliminary studies have shown that adding up to 12 cuts, on average, provides an optimal trade-off between lower bound improvements and an increase in computational complexity.
Instance Set and Instance Classes
Each of the aforementioned 5 configurations is used to solve the instance set generated as described in Section 6.1, amounting to a total of 1,215 instances per configuration. In total, 615 instances are infeasible. We omit said instances and compare the ascribed feature configurations on the remaining 600 instances. Appendix 15 contains additional details on the characteristics of all feasible instances.
In order to distinguish between easy and hard instances, we split the set of test instances into three categories. In preliminary studies, the worker strength has proven to have a significant impact on an instance’s complexity. Therefore, we consider instances with a worker strength between 0.3 and 0.5 as ‘hard’, while instances with a strength of 0.6 or 0.7 are seen as ‘medium’ and 0.8 or 0.9 as ‘easy’. In fact, almost all instances in the latter category were solved within a few seconds, whereas instances in the hard category are frequently not solved to optimality within the set time limit. We note that all instances with a worker strength of 0.2 or less are infeasible, while only 2 instances with a worker strength of 0.3 are feasible.
Unless otherwise stated, all tables in this section, contain average values. The number of instances on which the following analyses are based can be seen in Appendix 16.
Comparison to Benchmark Approaches
Table 4 compares the results obtained using the full configuration with the basic configuration.
Table 4: Comparison of Yuan, Liu, and Jiang (2015) with basic and full solver configuration
Complexity Config. % Opt Gap % UB LB Runtime
Easy Full 100.00% 0.00% 2.86 2.86 1.53s
Basic 82.76% 6.63% 2.86 2.12 33.30s
Yuan et al. 92.34% 2.12% 2.86 2.62 20.28s
Medium Full 90.43% 0.29% 15.54 15.48 31.84s
Basic 57.89% 7.48% 15.59 14.29 84.80s
Yuan et al. 63.46% 5.25% 15.54 14.56 80.45s
Hard Full 64.52% 2.77% 67.62 64.60 84.99s
Basic 49.61% 4.29% 69.19 64.84 106.08s
Yuan et al. 42.40% 4.95% 68.92 64.66 122.18s
All Full 89.23% 0.68% 20.84 20.19 29.61s
Basic 67.06% 6.44% 21.27 19.57 66.68s
Yuan et al. 71.72% 3.81% 21.20 19.86 62.79s
Column “% Opt” contains the percentage of instances that have been solved to optimality. It can be seen that the full configuration solves 22% more instances to optimality than the basic configuration. Furthermore, it returns significantly better average optimality gaps for all instance classes, where this effect decreases with increasing worker strength. While the full configuration returns substantially better lower bounds for all instance classes, the upper bounds also increased slightly for harder instances. Moreover, runtimes for small and medium instances reduce drastically, while larger instances are solved around 22 seconds faster on average. Additionally, it can be observed that, while the algorithm by Yuan, Liu, and Jiang (2015) is able to outperform the basic configuration for easy instances, it struggles to do so for medium and hard instances. Finally, the proposed full configuration manages to return the best runtimes, bounds, and optimality percentages out of all compared approaches for all complexity classes.
Solution Quality Robustness
Figure 3: Box plot of optimality gap percentage for all solver configurations
Figure 3 visualizes the frequencies of non-zero optimality gap percentages for all five considered solver configurations in a box plot. The red line is the median, while the boxes are limited by the 25%- and 75%-quantiles. The lower and upper whiskers are calculated as $q_{0.25}-1.5\cdot IQR$ and $q_{0.75}+1.5\cdot IQR$, respectively, where $IQR=q_{0.75}-q_{0.25}$ is the interquartile range. Clearly, the basic configuration has a high median optimality gap, as well as a comparable large set of outliers, sometimes even terminating with a gap of close to 70%. These metrics generally improve when enabling two of the three algorithm’s features, while the best results can be observed when using the full solver configuration. Additionally, larger gap percentages occur rather frequently when using the basic configuration, as all instances lie within their respective whiskers. Hence, the full solver configuration returns the best and most stable results from all compared strategies.
Full vs. no Branching
Table 5: Comparison of full and no branching on task finish times configuration
Complexity Config. % Opt Gap % UB LB Ropt Nopt Nexp Rnode Rpr
Easy Full 100.00% 0.00% 2.86 2.86 1.39s 1.28 1.54 0.99s 0.041s
No Branching 99.23% 0.27% 2.86 2.85 1.47s 1.93 6.50 0.45s 0.039s
Medium Full 90.43% 0.29% 15.54 15.48 5.79s 11.39 36.73 0.87s 0.111s
No Branching 73.68% 2.57% 15.58 15.04 7.00s 7.59 49.43 1.14s 0.219s
Hard Full 64.52% 2.77% 67.62 64.60 16.17s 21.00 65.24 1.30s 0.242s
No Branching 56.00% 3.73% 69.30 65.29 17.95s 17.95 50.32 1.90s 0.386s
All Full 89.23% 0.68% 20.84 20.19 4.74s 7.01 27.22 1.09s 0.108s
No Branching 81.18% 1.81% 21.29 20.25 5.31s 5.83 30.78 1.34s 0.175s
Table 5 compares the results obtained by the full solver configuration with the configuration ‘no branching’, i.e., branching on task finish times is disabled. Columns “Ropt” and “Nopt” describe the runtime and the number of explored nodes until optimality has been proven. We note that, unlike all other values in the above table, the baseline for these two metrics is not the entire instance set of a fixed instance class, but instead the set of instances that has been solved to optimality by both solver configurations. Column “Nexp” describes the number of explored nodes, while “Rnode” and “Rpr” contain the runtime per explored node and the runtime per pricing iteration, respectively. The average solution quality improves when the branching strategy is enabled, as 8.05% additional instances are solved to optimality, and the average gap is reduced by around 1.27%. Furthermore, solving a single node requires, on average, 25% less time, where this improvement increases for harder instances. This can be explained by significant time savings in the pricing step, where the runtime decreases from 0.386 seconds per iteration to 0.242 seconds.
For easy and medium instances, the full solver configuration requires less time to prove optimality and explores around 40% to 50% fewer nodes to do so, respectively. For the set of hard test instances, using branching on task finish times improves the solution quality noticeably, while optimality gaps are reduced by 1.04%. However, around 20% more nodes are required to prove optimality. This indicates that for larger instances, branching on task finish times is able to provide very good lower bounds, but does not excel at proving optimality. In summary, task finish times appear to be an efficient basis for branching decisions, as they significantly simplify the pricing step. For more complex instances, further studies on a potential tail-off effect on the quality of lower bounds and possible ways to mitigate this issue are needed.
Full vs. no DRMP
Table 6: Comparison of full and no DMP configuration
Complexity Config. % Opt Gap % UB LB Ropt Nopt ARMPn DRMPn NoAinf
Easy Full 100.00% 0.00% 2.86 2.86 1.44s 1.49 1.00 4.70 1.00
No DMP 100.00% 0.00% 2.86 2.86 1.50s 1.85 3.00 0.00 1.00
Medium Full 90.43% 0.29% 15.54 15.48 8.59s 18.63 44.50 7.90 1.12
No DMP 86.12% 0.57% 15.55 15.46 8.21s 17.57 88.80 0.00 19.44
Hard Full 64.52% 2.77% 67.62 64.60 24.18s 38.84 123.40 6.10 2.09
No DMP 63.71% 2.91% 67.55 64.42 20.14s 35.33 131.60 0.00 9.36
All Full 89.23% 0.68% 20.84 20.19 7.05s 12.48 63.40 7.20 1.38
No DMP 87.54% 0.81% 20.83 20.14 6.38s 11.82 94.30 0.00 15.18
[…]
Full vs. no CGCs
Table 7: Comparison of full and no Gomory cut configuration
Complexity Config. % Opt Gap % UB LB Rnode RneR % Root solved RootLB
Easy Full 100.00% 0.00% 2.86 2.86 0.99 0.71 92.72% 2.84
No CGCs 96.55% 0.46% 2.86 2.82 0.44 0.46 47.89% 1.97
Medium Full 90.43% 0.29% 15.54 15.48 0.87 0.76 50.72% 14.79
No CGCs 83.17% 1.22% 15.28 15.13 0.44 0.43 13.44% 13.13
Hard Full 64.52% 2.77% 67.62 64.60 1.30 1.00 29.03% 62.88
No CGCs 60.80% 2.85% 68.83 65.66 0.68 0.62 7.20% 62.26
All Full 89.23% 0.68% 20.84 20.19 1.09 0.88 64.65% 19.58
No CGCs 84.34% 1.23% 21.09 20.36 0.54 0.51 27.44% 18.56
Table 7 analyzes the results obtained by the full configuration and compares it to the ‘no CGCs’ configuration, i.e., no CGCs are added. Column “rootLB” contains the lower bound obtained after solving the root node, while column “RneR” depicts the algorithm’s runtime per node, excluding the root node. On average over all test instances, the usage of Gomory cuts improves the final lower bounds, allowing us to solve 4% more instances to optimality and closing the optimality gap by an additional 0.55%. Furthermore, the percentage of instances solved at the root node more than doubles from 27.44% to 64.65%, while it even increases four-fold for hard instances. Moreover, the lower bound at the root node improves by around 6%. Additionally, column “RnerR” shows that CGCs significantly increase the runtime per non-root node, doubling from 0.54 seconds to 1.09 seconds on average. Altogether, Gomory cuts aid to provide excellent bounds early on during the solving procedure. However, the increase in computational complexity, especially during the pricing step, implies a trade-off between bound quality and computational complexity. Especially for larger instances, carefully separating and electing CGCs plays a crucial role in the approach’s efficiency.
To conclude this section, we showed that our developed solution strategy is able to significantly improve the ‘basic’ configuration first described by Dall’Olio and Kolisch (2023), both with respect to optimality guarantees and bound quality. While the impact of every single configuration’s component seems to be rather small, combining our three core features and using them simultaneously greatly benefits the resulting algorithm’s performance.
6.3 Stochastic and Deterministic Solutions
In the following, we compare the quality of stochastic solutions with that of deterministic ones. For that purpose, all 1,215 test instances are solved using our stochastic approach and three deterministic approaches assuming best-case, median, or worst-case travel times. We note that we can adjust the AMP to use deterministic travel times $\tau_{i,j}$ for tasks $i$ and $j$ by assuming $P(t_{i,j}=\tau)=1$ when constructing the respective master problem. We thus obtained a total of four solutions for each instance, one for each type of deterministic travel time and one for stochastic travel times. We then sample 1,000 scenarios for each instance by randomly generating travel times for all pairs of parking positions according to the respective empirical distributions. For each instance and each scenario, we analyze the finish times of all tasks for all four solutions. By aggregating these results for each instance and solution over all 1,000 scenarios, we obtain an empirical service level and objective function value (10), which allows us to further assess the performance of deterministic and stochastic solutions. Furthermore, we use the deterministic framework, i.e., constraints (3) and (4) need to be satisfied. For deterministic travel times, this is equivalent to using hard time windows $\mathrm{ES}_{i},\mathrm{LF}_{i}]$ for each task $i$ and disallowing delays. In the following, we refer to the solution of an instance when using stochastic travel times as the stochastic solution of an instance. For best-, median, and worst-case travel times, we analogously refer to best-, median, and worst-case travel time solutions, respectively.
Table 8: Frequencies of stochastic (in)feasibility
Deterministic Stochastic
Complexity Travel Times #Feas. #alpha-feas. #LFv-feas. #Stoch.-feas.
Easy Best 266 11 29 11
Median 265 142 243 142
Worst 261 261 261 261
Stochastic 261 261 261 261
Medium Best 243 8 26 8
Median 228 105 192 103
Worst 203 203 203 203
Stochastic 209 209 209 209
Hard Best 222 0 17 0
Median 161 31 98 31
Worst 109 109 109 109
Stochastic 124 124 124 124
All Best 731 19 72 19
Median 654 278 533 276
Worst 573 573 573 573
Stochastic 594 594 594 594
Table 8 summarizes the feasibility of instances and solutions under various travel time assumptions. For each type of travel time, column “#Feas.” contains the number of instances whose master problem (10)–(13) is feasible, assuming the respective travel times. Furthermore, column “#Alpha-feas.” denotes the number of instances whose solutions satisfy the desired service level $\alpha$ for each task. This requirement is equivalent to satisfying constraint (3). Column “#LFe-feas.” serves a similar purpose and contains the number of instances whose solutions guarantee no delays beyond extended latest finish times. This corresponds to satisfying equality (4). Finally, column “#Stoch.-feas.” contains the number of instances whose solutions were feasible with respect to the prescribed requirements regarding both service levels and maximum delays. This is equivalent to fulfilling constraints (3) and (4) simultaneously. It is clear that a decrease in the available workforce increases the frequency with which the prescribed service level is not satisfied or large delays can not be ruled out. Furthermore, only 31 out of 728 best-case travel time solutions satisfy the desired service level for all tasks, while only 25 solutions are feasible with respect to both service level and maximum delay constraints. When median travel times are assumed, around 49% of all solutions guarantee the desired service level, while only 304 out of 649, i.e., around 47% of solutions also fulfill the maximum delay requirement (4). Thus, if deterministic travel times are assumed, it is likely that either tasks are delayed with a probability of more than $1-\alpha$ or large delays of more than 10 minutes can not be prevented, ultimately leading to frequent delays of multiple aircrafts.
Table 9: Objective function values and service levels
Objective Service Level
Complexity Travel Times Obj Obj-Pen $\mu_{SL}$ $\sigma_{SL}$ $\min_{SL}$
Easy Best 146.74 100.14 89.69% 7.99% 0.00%
Median 25.02 22.97 98.76% 1.32% 32.70%
Worst 3.21 3.21 100.00% 0.00% 100.00%
Stochastic 2.86 2.86 100.00% 0.03% 90.90%
Medium Best 146.11 107.25 91.39% 6.80% 0.20%
Median 33.51 31.86 98.90% 1.33% 28.50%
Worst 18.39 18.39 100.00% 0.00% 100.00%
Stochastic 15.68 15.65 99.96% 0.14% 90.20%
Hard Best 135.29 89.70 89.70% 9.50% 0.20%
Median 69.62 66.79 98.01% 1.81% 26.40%
Worst 68.43 68.43 100.00% 0.00% 100.00%
Stochastic 59.91 59.69 99.82% 0.18% 90.10%
All Best 154.09 109.35 90.29% 7.19% 0.20%
Median 36.14 34.45 98.67% 1.46% 26.40%
Worst 20.99 20.99 100.00% 0.00% 100.00%
Stochastic 18.25 18.23 99.96% 0.13% 90.10%
Table 9 provides an overview of objective function values and service levels. In order to have a stable foundation for comparison, all data visualized in Table 9 is calculated based on the set of instances for which feasible solutions are obtained when assuming worst-case travel times. This amounts to a total of 578 out of 1,215 instances. Recall that all values depicted here are calculated based on a Monte Carlo simulation of travel times. Columns “Obj” and “Obj-Pen” contain the average objective function value of deterministic solutions applied to stochastic travel times, with and without quadratic penalties for time window violations, respectively. Moreover, columns $\mu_{SL}$, $\sigma_{SL}$, and $\min_{SL}$ represent the average service level for all tasks, its standard deviation, and the minimum service level of any task. On average, stochastic solutions return the smallest objective function values, both with and without penalties, for all instance classes. While stochastic solutions’ objective function values are equal to around 3% of the objective function values of best-case travel time solutions (both excluding penalties, respectively) for easy instances, this relation reduces to around 35% for hard instances. Practically speaking, for easy instances, the total safety time buffer accumulated is 30 times larger when using stochastic rather instead of best-case travel time data. Moreover, for every instance class, stochastic solutions guarantee a very high and stable service level, averaging over 99.5% and having a standard deviation of 0.12%. Furthermore, while median travel time solutions provide relatively good objective function values, which are roughly 20% above stochastic solutions, they frequently provide very poor service level bounds for individual tasks, going as low as 21.8% for harder instances. Therefore, if a deterministic model is sought to solve the problem at hand, using median travel times appears to be the most promising approach. Nevertheless, service levels for individual tasks can still be arbitrarily low and clearly lack a lower bound. Such guarantees can only be made with stochastic solutions, which outperform deterministic travel time solutions by all previously mentioned metrics.
Figure 4: Distribution of service levels
Figure 4 shows the distribution of service levels per task. For this purpose, for each type of travel time, we collected all service levels for each individual task for every instance and plotted their empirical distribution. While stochastic travel times guarantee a service level of at least 90% at all times, median and best-case travel time solutions frequently provide significantly lower values, down to 3% in some cases. Moreover, best-case travel time solutions violate the prescribed minimum service level for 25% of tasks, while for median travel time solutions, this holds in around 8% of cases. This emphasizes the unpredictability of minimum service levels when deterministic travel times are assumed, which ultimately has a significant impact on perceived service quality. Hence, explicitly considering stochastic travel times in the problem formulation yields solutions that seldom violate any time windows, resulting in a high and reliable quality of service.
Figure 5: Frequency of durations of time window violations
Figure 5 visualizes the occurrence of time window violations, measured in time steps. Recall that for our purposes, each time step is 2 minutes long. As for Figure 4, we collected all potential time window violations and their lengths for each instance and each scenario. Overall, best-case travel time solutions have higher median time window violations, larger 75%-quantiles, and more outliers than stochastic solutions. At the same time, the median time window violations and 75%-quantiles of median and stochastic solutions are almost identical. Nevertheless, the former solutions exhibit far more large delays, reaching up to 11 time steps, i.e., 22 minutes. In total, assuming deterministic travel times usually returns solutions that cause comparably long delays, are at risk of delaying tasks by significant amounts, and lead to highly volatile delays, which are several magnitudes larger than the maximum acceptable delay $\mathrm{LF}_{i}^{\mathrm{e}}-\mathrm{LF}_{i}$. These effects lead to passengers becoming dissatisfied and baggage handling operators having to pay financial penalties. Using stochastic travel times greatly benefits a solution’s quality in these regards, typically causing few, small delays and ruling out undesirably long delays.
To summarize the above findings, we observe that solutions based on deterministic travel times, be they best-case, median, or worst-case, return strategies that do not live up to the set service quality. Explicitly incorporating stochastic travel times into the model allows for solutions that efficiently use the available workforce to increase safety time buffers, guarantee a stable service level, and prevent large delays.
7 Conclusion
We have examined the problem of forming and routing worker teams, consisting of workers of different hierarchical skill levels. We extended the works of Dall’Olio and Kolisch (2023) by considering stochastic travel times across the apron. For this purpose, we adjusted their MIP formulation to suit stochastic travel times, provide a service level guarantee, and disallow extreme delays for each task. Furthermore, we introduced an alternative master problem formulation abbreviated as DRMP. Moreover, we provided an adjusted dominance rule used to solve the pricing problem associated with said model, allowing us to solve multiple pricing networks at once, at the expense of only a slight increase in runtime. This result ultimately made the usage of said formulation practically feasible. We then combined the DRMP with the formulation proposed by the aforementioned authors to develop a new solution approach based on a Branch-Price-and-Cut approach. More specifically, we switch between two master problem formulations, depending on the characteristics of the current (integer) solution and the position within the branching tree. Furthermore, we introduced an improved branching strategy based on the worst-case finish times of individual tasks, which was able to greatly improve the solution quality and algorithm’s convergence speed. Moreover, we applied an exact separation procedure for rank-1 Chvátal-Gomory cuts and carefully designed the decision process of when and how to separate said cuts. Our computational studies indicate that our proposed algorithm is able to outperform the standard Branch-Price-and-Check approach by Dall’Olio and Kolisch (2023), solving more instances to optimality and significantly reducing optimality gaps. Additionally, we observed that including stochastic travel times into the model is a superior approach to assuming deterministic travel times, as the solutions obtained allow for earlier expected termination of tasks, less frequent and smaller delays, and a guaranteed, stable service level and maximum possible delay.
Our computational studies have shown that the impact of branching on task finish times on upper and lower bounds gradually decreases with increasing problem size. Hence, it might be worthwhile to further analyze the reasons for this phenomenon and identify more advanced branching strategies on task time windows that scale well with the problem size. Moreover, one of our core assumptions is the separation of daily flight schedules into segments of 1 to 2 hours, within which optimal decisions can be made without considering what happens before or after this. While this is a realistic assumption for Munich Airport, it might not hold for other airports. Similarly, if longer time horizons are considered, time windows might also be subject to significant uncertainty. In order to derive reasonable team formation and routing decisions for such settings, one might consider an infinite time horizon, potentially applying multi-stage stochastic programming or online optimization techniques. Furthermore, as we consider task execution times and time windows to be deterministic, our model is still a simplification. As the exact solvability of a full-fledged stochastic model that incorporates all potential factors of stochasticity is questionable, further advances toward heuristic approaches can be made. For instance, the order in which the pricing networks are solved in each pricing step has a large impact on the algorithm’s runtime. Moreover, pricing networks can be solved heuristically, potentially returning a very good, yet not optimal column, in a fraction of the runtime needed by an exact labeling approach. Such decision processes could be further improved by applying machine learning techniques tailored individually to the problem’s structure. Finally, the applicability of the aggregation technique of multiple pricing networks introduced in Section 5.3 to other problem types, such as different variants of the VRP, where the SPPRC arises as a pricing subproblem and several subproblems share many of their properties, can be further studied. In particular, multi-depot VRPs and VRPs with mixed fleets can be of special interest for this.
[…]
We thank Dr. Jan Evler and Aeroground GmbH for providing the data and key industry information. Andreas Hagn and Giacomo Dall’Olio have been funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) - Project number 277991500.